### v1

# Initial-Condition-Robust Inference in Autoregressive Models

Abstract
This paper considers confidence intervals (CIs) for the autoregressive (AR) parameter in an AR model with an AR parameter that may be close or equal to one. Existing CIs rely on the assumption of a stationary or fixed initial condition to obtain correct asymptotic coverage and good finite sample coverage. When this assumption fails, their coverage can be quite poor. In this paper, we introduce a new CI for the AR parameter whose coverage probability is completely robust to the initial condition, both asymptotically and in finite samples. This CI pays only a small price in terms of its length when the initial condition is stationary or fixed. The new CI also is robust to conditional heteroskedasticity of the errors.
Keywords: Asymptotic size, autoregressive model, confidence set, initial condition, robustness.
JEL Classification Numbers: C10, C12.
1.  Introduction
We consider an AR model with an AR parameter that may take values close or equal to one. Existing CIs in the literature assume that the initial condition is stationary, zero, or fixed, e.g., see [14], [7], [1], [12], [9], [13], and [4, AG14 hereafter]. If the initial condition assumption is not satisfied, then these CIs do not have correct asymptotic coverage probabilities and their finite sample coverage probabilities can be poor. For example, simulations reported below show that the nominal 95% AG14 CI has finite sample coverage probabilities ranging from 24.1% to 93.5% across 50 cases with highly variable initial conditions and different types of conditional heteroskedasticity of the error. A quarter of these CPs are 79.0% or less.
To circumvent the sensitivity of existing CIs to the initial condition, this paper introduces a new initial-condition-robust (ICR) CI whose finite sample CPs do not depend on the initial condition. These CIs are easy to compute and do not depend on any tuning parameters. We show that the ICR CI has correct asymptotic size in a uniform sense under conditions that allow for arbitrary initial condition and conditional heteroskedasticity. Simulations show that the finite sample CPs of the nominal 95% ICR CI are quite good. They lie between 93.5% and 95.0% in the cases considered.
Like the other CIs referenced above, the ICR CI is constructed by inverting hypothesis tests concerning the value of the AR parameter. For all of these CIs, the tests are based on a t statistic that depends on a least squares (LS) estimator of the AR parameter. In contrast to existing CIs, the LS estimator used by the ICR tests contains an additional regressor that eliminates the effect of the initial condition under the null hypothesis. (For the definition of this regressor, see Section 2 below.) In consequence, the ICR CI has a CP that is invariant to the value of the initial condition. However, the length of the ICR CI can be effected by the initial condition. Simulations show that the ICR CI is slightly shorter when the initial condition is highly variable than when it is stationary or fixed. [^1] The length of the ICR CI varies mostly with the value of the AR parameter. It is shortest for AR parameters near one. This is true of existing CIs as well.
In scenarios where existing CIs are asymptotically valid (i.e., the initial condition is stationary or fixed), the ICR CI pays a price in terms of its length compared to existing CIs. This occurs because the ICR LS estimator has a higher variance due to the additional regressor that is included in the LS regression. However, simulations show that the effect is fairly small. Across 50 scenarios with stationary or fixed initial conditions and i.i.d., GARCH, or ARCH errors, the ratio of the expected length of the ICR CI compared to that of the AG14 CI was found to vary between 1.00 and 1.11. On average, the ICR CIs are 3.5% longer than the AG14 CIs in these scenarios.
Based on the ICR CI, we also introduce an asymptotically median-unbiased interval estimator (MUE) of the AR parameter.
This paper is organized as follows. Section 2 specifies the AR model and defines the ICR CI. Section 3 defines the median-unbiased interval estimator. Section 4 provides simulation results concerning CPs and average lengths of the ICR and AG14 CIs and absolute median biases of the ICR MUE. Section 5 provides asymptotic results for the ICR CI. The Supplemental Material describes how the critical values were computed, proves Theorems 1 and 2, and provides some additional simulation results.
2.  Model and ICR Confidence Interval
We consider the AR(1) model with conditional-heteroskedasticity studied in AG14:
$Y_{i}$ $=\mu+Y_{i}^{*}\quad\text{and}$
$Y_{i}^{*}$ $=\rho Y_{i-1}^{*}+U_{i}\quad\text{for }i=1,2,\dots,n,$ (1)
where $\rho\in[-1+\varepsilon,1]$ for some $0<\varepsilon<2$ and $\{U_{i}:i=1,2,\dots\}$ are stationary and ergodic under the distribution $F$, with conditional mean 0 given a $\sigma$-field $\mathcal{G}_{i-1}$ for which $U_{j}\in\mathcal{G}_{i}$ for all $j\leq i$, conditional variance $\sigma_{i}^{2}=E_{F}(U_{i}^{2}|\mathcal{G}_{i-1})$, and unconditional variance $\sigma_{U}^{2}\in(0,\infty)$. A detailed description of the parameter space $\Lambda$ is provided in Section 5.1.
By iterative substitution in (1), we have
$Y_{i}$ $=\mu+\rho^{i-1}\cdot\rho\cdot Y_{0}^{*}+\sum_{j=1}^{n}\rho^{i-j}U_{j}\quad\text{for }i=1,\dots,n.$ (2)
Let $Y$, $U$, and $X_{1}$ be $n$-vectors whose $i$-th elements are $Y_{i}$, $U_{i}$, and $Y_{i-1}$, respectively. Let $X_{2}(\rho)$ be an $n\times 2$ matrix whose $i$-th row is $(1,\rho^{i-1})$ when $1<\rho<1$. [^2] Following the convention $0^{0}=1$, we use $\rho^{i-1}$ in the regression instead of $\rho^{i}$ to avoid perfect multicollinearity when $\rho=0$. When $\rho=1$, we define $X_{2}(\rho)$ to be an $n\times 2$ matrix whose $i$-th row is $(1,i)$. Let $X(\rho)=[X_{1}:X_{2}(\rho)]$, $P_{X(\rho)}=X(\rho)(X(\rho)^{\prime}X(\rho))^{-1}X(\rho)^{\prime}$, and $M_{X(\rho)}=I_{n}-P_{X(\rho)}$. Let $\widehat{U}_{i}$ denote the $i$-th element of the residual vector $M_{X(\rho)}Y$. Let $p_{ii}$ denote the $i$-th diagonal element of $P_{X(\rho)}$, and define $p_{ii}^{*}=\min\{p_{ii},n^{-1/2}\}$. Let $\Delta$ be a diagonal $n\times n$ matrix whose $i$-th diagonal element is $\widehat{U}_{i}/(1-p_{ii}^{*})$. Then, the ICR LS estimator $\widehat{\rho}_{n}(\rho)$, and the HC5 variance estimator $\widehat{\sigma}_{n}^{2}(\rho)$ are defined as
$\widehat{\rho}_{n}(\rho)=$ $\ (X_{1}^{\prime}M_{X_{2}(\rho)}X_{1})^{-1}X_{1}^{\prime}M_{X_{2}(\rho)}Y\text{ and}$
$\widehat{\sigma}_{n}^{2}(\rho)=$ $\ (n^{-1}X_{1}^{\prime}M_{X_{2}(\rho)}X_{1})^{-1}(n^{-1}X_{1}^{\prime}M_{X_{2}(\rho)}\Delta^{2}M_{X_{2}(\rho)}X_{1})(n^{-1}X_{1}^{\prime}M_{X_{2}(\rho)}X_{1})^{-1},$ (3)
respectively.
Let
$T_{n}(\rho)=\frac{n^{1/2}(\widehat{\rho}_{n}(\rho)-\rho)}{\widehat{\sigma}_{n}^{2}(\rho)}.$ (4)
For suitable sequences $\{\rho_{n}\}_{n\geq 1}$ such that $n(1-\rho_{n})\to h\in[0,\infty]$, we show that
$T_{n}(\rho_{n})\stackrel{{\scriptstyle d}}{{\to}}J_{h},$ (5)
where $J_{h}$ is defined in Section 5.2 below. Table 1 provides the quantiles $c_{h}(\alpha/2)$ and $c_{h}(1-\alpha/2)$ of the distribution of $J_{h}$ for $\alpha=.05$ and .1.1, which are the critical values employed by the ICR CI.
The nominal $1-\alpha$ equal-tailed two-sided ICR CI for $\rho$ is
$CI_{\text{ICR},n}:=\left\{\rho\in[-1+\epsilon,1]:c_{h}(\alpha/2)\leq T_{n}(\rho)\leq c_{h}(1-\alpha/2)\ \text{for}\ h=n(1-\rho)\right\},$ (6)
which can be computed by taking a fine grid of values $\rho\in[-1+\epsilon,1]$ and comparing $T_{n}(\rho)$ to $c_{h}(\alpha/2)$ and $c_{h}(1-\alpha/2)$. Given these critical values, computation of the CIs is fast.
Table 1: Quantiles of $J_{h}$ for Use with 90% and 95% Equal-Tailed Two-Sided CIs and MUEs
Values of $c_{h}\left(\alpha\right)$, the $\alpha^{\text{th}}$ Quantile of $J_{h}$, for Use with 90% and 95% Equal-Tailed Two-Sided CI’s and MUE’s
$h$ 0 .2 .4 .6 .8 1 1.4 1.8 2.2 2.6 3 3.4 3.8
$c_{h}\left(.025\right)$ -3.66 -3.63 -3.60 -3.56 -3.54 -3.52 -3.46 -3.40 -3.36 -3.31 -3.27 -3.23 -3.19
$c_{h}\left(.05\right)$ -3.41 -3.38 -3.35 -3.31 -3.28 -3.25 -3.20 -3.14 -3.08 -3.04 -3.00 -2.95 -2.91
$c_{h}\left(.5\right)$ -2.18 -2.13 -2.09 -2.04 -1.99 -1.95 -1.86 -1.78 -1.70 -1.63 -1.57 -1.50 -1.45
$c_{h}\left(.95\right)$ -.94 -.87 -.80 -.74 -.68 -.62 -.50 -.39 -.29 -.19 -.11 -.03 .05
$c_{h}\left(.975\right)$ -.65 -.59 -.52 -.45 -.38 -.32 -.21 -.08 .01 .11 .19 .28 .35
$h$ 4.2 4.6 5 6 7 8 9 10 11 12 13 14 15
$c_{h}\left(.025\right)$ -3.16 -3.12 -3.09 -3.02 -2.97 -2.90 -2.87 -2.82 -2.79 -2.75 -2.73 -2.71 -2.69
$c_{h}\left(.05\right)$ -2.87 -2.83 -2.80 -2.72 -2.66 -2.61 -2.56 -2.52 -2.48 -2.45 -2.42 -2.40 -2.38
$c_{h}\left(.5\right)$ -1.39 -1.34 -1.30 -1.20 -1.11 -1.04 -.99 -.93 -.89 -.85 -.82 -.78 -.76
$c_{h}\left(.95\right)$ .11 .18 .24 .36 .46 .55 .61 .68 .74 .78 .81 .84 .88
$c_{h}\left(.975\right)$ .41 .48 .54 .66 .77 .86 .92 .99 1.05 1.08 1.12 1.15 1.20
$h$ 20 25 30 40 50 60 70 80 90 100 200 300 500
$c_{h}\left(.025\right)$ -2.59 -2.53 -2.47 -2.41 -2.36 -2.32 -2.30 -2.27 -2.26 -2.25 -2.16 -2.13 -2.09
$c_{h}\left(.05\right)$ -2.28 -2.22 -2.15 -2.09 -2.05 -2.01 -1.99 -1.96 -1.94 -1.94 -1.84 -1.81 -1.78
$c_{h}\left(.5\right)$ -.65 -.58 -.52 -.45 -.41 -.37 -.34 -.32 -.30 -.28 -.20 -.16 -.13
$c_{h}\left(.95\right)$ .99 1.06 1.12 1.19 1.24 1.28 1.30 1.32 1.34 1.36 1.45 1.48 1.52
$c_{h}\left(.975\right)$ 1.30 1.38 1.43 1.50 1.55 1.59 1.62 1.64 1.66 1.68 1.76 1.79 1.83
3.  ICR Median-Unbiased Interval Estimator
By definition, an estimator $\widehat{\theta}_{n}$ of a parameter $\theta$ is median unbiased if $P(\widehat{\theta}_{n}\geq\theta)\geq 1/2$ and $P(\widehat{\theta}_{n}\leq\theta)\geq 1/2$. In this section, we introduce an ICR MUE of $\rho$ that satisfies an analogous condition. Also, with probability close to one, this interval estimator is a point. [^3] If the median of the asymptotic null distribution of the $t$-statistic is strictly decreasing in $\rho$, or equivalently, if $c_{h}(0.5)$ is strictly increasing in $h$, then the proposed interval estimator would be a point estimator with probability one. Because this condition fails to hold exactly, but almost holds, there is a very small probability that the estimator is a short interval rather than a point.
Let $CI_{\text{ICR},n}^{\text{up}}(.5)$ and $CI_{\text{ICR},n}^{\text{low}}(.5)$ denote level .5.5 one-sided upper-bound and lower-bound CIs for $\rho$, respectively. By definition,
$CI_{\text{ICR},n}^{\text{up}}(.5)$ $\coloneqq\left\{\rho\in[-1+\epsilon,1]:c_{h}(0.5)\leq T_{n}(\rho)\ \text{for}\ h=n(1-\rho)\right\}\text{ and}$
$CI_{\text{ICR},n}^{\text{low}}(.5)$ $\coloneqq\left\{\rho\in[-1+\epsilon,1]:T_{n}(\rho)\leq c_{h}(0.5)\ \text{for}\ h=n(1-\rho)\right\}.$ (7)
The MUE $\widetilde{\rho}_{n}$ of $\rho$ is defined by
$\widetilde{\rho}_{n}$ $=[\widetilde{\rho}_{n}^{\text{low}},\widetilde{\rho}_{n}^{\text{up}}],\text{ where}$
$\widetilde{\rho}_{n}^{\text{up}}$ $=\max\{\rho:\rho\in CI_{\text{ICR},n}^{\text{up}}(.5)\}\text{ and}$
$\widetilde{\rho}_{n}^{\text{low}}$ $=\min\{\rho:\rho\in CI_{\text{ICR},n}^{\text{low}}(.5)\}.$ (8)
By construction, we have $\widetilde{\rho}_{n}^{\text{low}}\leq\widetilde{\rho}_{n}^{\text{up}}$. [^4] This holds because $\widetilde{\rho}_{n}^{\text{up}}\geq\sup\{\rho\in[-1+\epsilon,1]:c_{h}(.5)=T_{n}(\rho)\}$ and $\widetilde{\rho}_{n}^{\text{low}}$ is less than or equal to the infimum of the values in the same set. In addition, $\widetilde{\rho}_{n}$ is a singleton whenever the set $\{\rho\in[-1+\epsilon,1]:T_{n}(\rho)=c_{h}(.5)\}$ contains a single point, in which case $\widetilde{\rho}_{n}$ equals this point. Table 1 provides the critical values $c_{h}(.5)$ for a wide range of $h$. Given these critical values, computation of $\widetilde{\rho}_{n}$ is fast.
4.  Monte Carlo Simulations
We compare the coverage probabilities (CPs) and average lengths (ALs) of the ICR and AG14 CIs and report the absolute median biases (ABMs) of the ICR MUE using Monte Carlo simulations. In Section D of the Supplemental Material, we compare the performance of the ICR and [13, Mik07 hereafter] CIs. We focus on the nominal 95% equal-tailed two-sided CIs. For the AG14 CI, we follow the paper’s constructions of the test statistic and critical values. To calculate the MUE for $\rho$, we follow the procedure described in Section 3 and use $\widetilde{\rho}_{n}$ as the MUE of $\rho$ when $\widetilde{\rho}_{n}$ is a singleton. When $\widetilde{\rho}_{n}^{\text{up}}\neq\widetilde{\rho}_{n}^{\text{low}}$, we take $\widetilde{\rho}_{n}^{\text{up}}$ as the MUE for $\rho$ following [5, Section 5.1].
We consider a wide range of $\rho$ values: 0, 0.5, 0.7, 0.9, and 0.99. The innovations are of the form $U_{i}=\sigma_{i}\varepsilon_{i}$, where $\{\varepsilon_{i}:i=0,\pm 1,\dots\}$ are identically and independently distributed (i.i.d.) standard normal and $\sigma_{i}$ is the multiplicative conditional heteroskedasticity. Let GARCH-$(ma,ar;\psi)$ denote a GARCH(1, 1) process with MA, AR, and intercept parameters $(ma,ar;\psi)$, and let ARCH-$(ar_{1},\dots,ar_{4};\psi)$ denote an ARCH(4) process with AR parameters $(ar_{1},\dots,ar_{4})$ and intercept $\psi$. Following AG14, we consider five specifications for the conditional heteroskedasticity of the innovations: (a) i.i.d. $N(0,1)$ (“i.i.d.”), (b) GARCH(1,1)-(.05,.9;.001) (“GARCH1”), (c) GARCH(1,1)-(.15,.8;.2) (“GARCH2”), (d) GARCH(1,1)-(.25,.7;.2) (“GARCH3”), and (e) ARCH(4)-(.3,.2,.2,.2;.2) (“ARCH”). For the initial conditions, we consider the following four specifications: (a) Fixed: $Y_{0}^{*}=0$, (b) Stationary: $Y_{0}^{*}=\sum_{i=0}^{\infty}\rho^{i}U_{-i}$, (c) Scaled $n$: $Y_{0}^{*}\sim\sqrt{n}\cdot\sum_{i=0}^{\infty}\rho^{i}U_{-i}$, and (d) Explosive: $Y_{0}^{*}\sim n^{3/4}\cdot\sum_{i=0}^{\infty}\rho^{i}U_{-i}$. We consider a sample size of $n=150$. All simulation results are based on 30,000 simulation repetitions.
Tables 2 and 3 report the CPs of AG14 and ICR CIs, respectively. The CPs of the AG14 CI in Table 2 are close to the nominal level of 95% when the initial condition is fixed at zero or stationary, but fall below 95% when the initial condition is more variable, namely in the scaled $n$ and explosive $Y_{0}^{*}$ cases. In particular, the AG14 CI is robust to non-i.i.d. innovations when the initial condition is fixed or stationary. On the other hand, for scaled $n$ and explosive $Y_{0}^{*}$ cases, the AG14 CPs lie in the interval [24.1, 93.5] with roughly a quarter of the CPs being less than 79.0. Table 3 shows that the ICR CI achieves CPs close to 95% in all of the scenarios considered and for any initial conditions.
Table 4 reports the ratios of the ALs of the nominal 95% ICR CI to the AG14 CI. We include only cases with fixed or stationary initial conditions because AG14 CIs are not robust to more variable initial conditions, whereas ICR CIs are. The ratios are between 1.00 and 1.11. On average, the ICR CIs are 3.5% longer than the AG14 CIs over the fifty cases considered. Thus, the price the ICR CI pays in terms of AL to gain robustness against the distribution of the initial condition is fairly small.
Table 2: Coverage probabilities ($\times 100$) of nominal 95% AG14 CIs
Initial Conditions: Fixed Stationary
$\bm{\rho}$: .00 .50 .70 .90 .99 .00 .50 .70 .90 .99
i.i.d. 94.6 94.6 94.8 94.7 95.1 94.6 94.5 94.8 94.7 94.5
GARCH1 94.6 94.5 94.8 94.8 95.1 94.6 94.6 94.9 94.8 94.4
GARCH2 94.4 94.6 94.5 94.9 95.2 94.3 94.6 94.5 94.8 94.7
GARCH3 94.2 94.3 94.6 94.7 95.2 94.1 94.3 94.6 94.6 94.6
ARCH4 93.7 93.9 94.2 94.6 95.6 93.6 94.0 94.1 94.5 94.7
Initial Conditions: Scaled $\bm{n}$ Explosive
i.i.d. 88.4 90.5 91.8 92.7 83.4 63.7 80.5 86.6 90.8 77.7
GARCH1 59.1 79.0 86.4 91.7 83.2 24.1 69.9 82.3 90.0 77.3
GARCH2 90.2 91.8 92.4 93.0 83.9 71.0 83.3 87.8 90.9 77.4
GARCH3 90.9 92.1 92.5 93.4 84.2 74.4 84.9 88.4 91.4 77.7
ARCH4 91.1 92.3 92.8 93.5 84.4 76.7 86.3 89.4 91.8 77.9
Table 3: Coverage probabilities ($\times 100$) of nominal 95% ICR CIs
Initial Conditions: Arbitrary
$\bm{\rho}$: .00 .50 .70 .90 .99
i.i.d. 94.4 94.5 94.7 94.7 94.3
GARCH1 94.4 94.6 94.9 95.0 94.3
GARCH2 94.1 94.6 94.4 94.9 94.2
GARCH3 93.9 94.2 94.5 94.7 94.1
ARCH4 93.5 93.8 93.9 94.5 94.3
Table 4: Ratios of average lengths of nominal 95% ICR to AG14 CIs
Initial Conditions: Fixed Stationary
$\bm{\rho}$: .00 .50 .70 .90 .99 .00 .50 .70 .90 .99
i.i.d. 1.00 1.00 1.02 1.04 1.08 1.00 1.01 1.02 1.05 1.11
GARCH1 1.04 1.05 1.06 1.07 1.08 1.00 1.01 1.02 1.05 1.11
GARCH2 1.00 1.01 1.03 1.02 1.07 1.00 1.01 1.03 1.04 1.11
GARCH3 1.00 1.01 1.03 1.02 1.06 1.00 1.02 1.04 1.03 1.10
ARCH4 1.00 1.02 1.03 1.01 1.06 1.01 1.02 1.04 1.03 1.10
Table 5 presents the ALs of ICR CIs. These are substantially decreasing in $\rho$, as expected, longer for ARCH than i.i.d., and increasing from GARCH1 to GARCH2 to GARCH3.
Table 5: Average lengths of nominal 95% ICR CIs
Initial Conditions: Fixed Stationary
$\bm{\rho}$: .00 .50 .70 .90 .99 .00 .50 .70 .90 .99
i.i.d. .32 .28 .24 .17 .08 .32 .28 .24 .17 .07
GARCH1 .33 .30 .25 .17 .08 .34 .30 .26 .17 .08
GARCH2 .38 .34 .29 .19 .08 .38 .34 .29 .19 .08
GARCH3 .43 .38 .33 .20 .09 .43 .38 .33 .20 .08
ARCH4 .49 .44 .37 .21 .09 .49 .44 .37 .21 .09
Initial Conditions: Scaled $\bm{n}$ Explosive
i.i.d. .31 .27 .23 .14 .04 .28 .23 .18 .09 .02
GARCH1 .28 .24 .19 .12 .04 .18 .15 .12 .07 .02
GARCH2 .37 .32 .26 .15 .04 .32 .27 .21 .10 .02
GARCH3 .41 .36 .29 .16 .04 .36 .30 .23 .11 .02
ARCH4 .46 .40 .32 .17 .04 .39 .32 .25 .12 .02
Table 6 reports the AMBs of the ICR MUE. Across all cases, the AMBs range from 0.000 to 0.022, indicating that their magnitudes are generally small.
Table 6: Absolute Median Biases of ICR MUEs
Initial Conditions: Fixed Stationary
$\bm{\rho}$: .00 .50 .70 .90 .99 .00 .50 .70 .90 .99
i.i.d. .012 .011 .006 .005 .020 .012 .011 .006 .005 .015
GARCH1 .017 .011 .006 .005 .020 .017 .011 .006 .005 .015
GARCH2 .017 .011 .006 .000 .020 .017 .011 .006 .000 .015
GARCH3 .022 .011 .011 .000 .020 .022 .011 .011 .000 .015
ARCH4 .022 .016 .011 .000 .020 .022 .016 .011 .000 .015
Initial Conditions: Scaled $\bm{n}$ Explosive
i.i.d. .012 .011 .006 .005 .010 .012 .011 .006 .010 .010
GARCH1 .017 .011 .006 .005 .010 .017 .011 .011 .010 .010
GARCH2 .017 .011 .011 .005 .010 .017 .011 .011 .010 .010
GARCH3 .022 .011 .011 .005 .010 .022 .011 .011 .010 .010
ARCH4 .022 .016 .011 .005 .010 .022 .016 .011 .010 .010
In Section D of the Supplemental Material, we compare the performance of the ICR and Mik07 CIs via simulations. The results are similar to those reported above for the ICR and AG14 CIs except that the Mik07 CI undercovers both under conditional heteroskedasticity, and/or scaled $n$ or explosive initial conditions, but by a lower magnitude in the latter case than the AG14 CI. Again, we find that the ICR CI is robust to more variable initial conditions while paying only a small price in terms of CI width in the situations where the Mik07 CI has correct asymptotic coverage, i.e., when the innovations are i.i.d. and the initial conditions are stationary (or fixed).
5.  Asymptotic Results
This section establishes the correct uniform asymptotic size and asymptotic similarity of the ICR CI for $\rho$.
5.1.  Parameter Space, Correct Uniform Size, and Asymptotic
Similarity
Following AG14, the parameter space for $(\rho,F)$ is given by:
$\Lambda=\Big\{\lambda=(\rho,F)$: (i) $\rho\in[-1+\epsilon,1]$;
  1. (ii)
    $\{U_{i}:i=1,2,\dots\}$ are stationary and strong-mixing under $F$ with $E_{F}(U_{i}|\mathcal{G}_{i-1})=0$ a.s., $E_{F}(U_{i}^{2}|\mathcal{G}_{i-1})=\sigma_{i}^{2}$ a.s., where $\mathcal{G}_{i}$ is some non-decreasing sequence of $\sigma$-fields for which $U_{j}\in\mathcal{G}_{i}$ for all $j\leq i$ for $i=1,2,\dots$;
  2. (iii)
    The strong-mixing numbers $\{\alpha_{F}(m):m\geq 1\}$ satisfy $\alpha_{F}(m)\leq Cm^{-3\zeta/(\zeta-3)}$, $\forall m\geq 1$;
  3. (iv)
    $\sup_{i,s,t}E_{F}|\prod_{a\in A}a|^{\zeta}\leq M$, where $0<i,s,t<\infty$, $i\geq\max(s,t)$, and $A$ is any non-empty subset of $\{U_{i-s},U_{i-t},U_{i+1}^{2},U_{1}^{2}\}$;
  4. (v)
    $E_{F}U_{1}^{2}=1$;
for some constants $0<\epsilon<2$, $\zeta>3$, and $C<\infty$$\Big\}$.
We suppose $E_{F}U_{1}^{2}=1$ for simplicity because the asymptotic distribution of $T_{n}(\rho)$ does not depend on the unconditional variance. Parts (ii) and (iv) are slightly different from the corresponding conditions in AG14 because the random variables $\{U_{-j},\ j\geq 0\}$ do not enter into the asymptotic analysis of the ICR CI. The key innovation of this paper is that we allow arbitrary initial conditions $Y_{0}^{*}$, including those corresponding to explosive processes.
The main theoretical result of this paper shows that $CI_{\text{ICR},n}$ has correct asymptotic size for the parameter space $\Lambda$ and is asymptotically similar. Let $P_{\lambda}$ denote probability under $\lambda=\left(\rho,F\right)\in\Lambda$.
Theorem 1.
Let $\alpha\in(0,1)$. For the parameter space $\Lambda$, the nominal $1-\alpha$ ICR CI for the parameter $\rho$ satisfies
$\underset{n\to\infty}{\lim\inf}\inf_{\lambda\in\Lambda}P_{\lambda}(\rho\in CI_{\text{ICR},n})=\underset{n\to\infty}{\lim\sup}\sup_{\lambda\in\Lambda}P_{\lambda}(\rho\in CI_{\text{ICR},n})=1-\alpha.$
The MUE $\widetilde{\rho}_{n}$ has the following median-unbiasedness property. This result is a corollary to the one-sided versions of Theorem 1.555Corollary 1 holds because $CI_{\text{ICR},n}^{\text{up}}(.5)$ and $CI_{\text{ICR},n}^{\text{low}}(.5)$ both have coverage probabilities of $1/2$ or greater by the proof of Theorem 1 applied to these one-sided CIs and $(\widetilde{\rho}_{n}\geq\rho)\supset CI_{\text{ICR},n}^{\text{up}}(.5)$ and $(\widetilde{\rho}_{n}\leq\rho)\supset CI_{\text{ICR},n}^{\text{low}}(.5)$.
Corollary 1.
The MUE estimator $\widetilde{\rho}_{n}$ satisfies
$\underset{n\to\infty}{\lim\inf}\inf_{\lambda\in\Lambda}P_{\lambda}(\widetilde{\rho}_{n}^{\text{up}}\geq\rho)$ $\geq 1/2\text{ and }\underset{n\to\infty}{\lim\inf}\inf_{\lambda\in\Lambda}P_{\lambda}(\widetilde{\rho}_{n}^{\text{low}}\leq\rho)\geq 1/2.$
5.2.  Asymptotic Distribution
As stated in (5), $J_{h}$ is the asymptotic distribution of $T_{n}(\rho_{n})$ under suitable sequences $\{\rho_{n}:n\geq 1\}$. Here, we state this result formally as Theorem 2 and define $J_{h}$. This result is proved in a similar way to Theorem 1 in [3]. Using Theorem 2.1 in [2], Theorem 2 is sufficient to establish Theorem 1, which is proved in Section B of the Supplemental Material.
Theorem 2.
For a sequence $\lambda_{n}=(\rho_{n},F_{n})$ such that $n(1-\rho_{n})\to h\in[0,\infty]$, we have
$T_{n}(\rho_{n})=\frac{n^{1/2}(\widehat{\rho}_{n}(\rho_{n})-\rho_{n})}{\widehat{\sigma}_{n}(\rho_{n})}\stackrel{{\scriptstyle d}}{{\to}}J_{h},$
where $J_{h}$ is defined immediately below.
Remark 1.
For any subsequence $\{p_{n}\}_{n\geq 1}$ of $\{n\}_{n\geq 1}$, Theorem 2 holds with $p_{n}$ in place of $n$ throughout and $h_{p_{n}}$ in place of $h=h_{n}$.
For $h=\infty$, $J_{h}$ is the $N(0,1)$ distribution. For $h\in[0,\infty)$,
$J_{h}:=\int_{0}^{1}I_{f,h}(r)dW(r)\Big/\left(\int_{0}^{1}I_{f,h}(r)^{2}dr\right)^{1/2},$ (9)
where $I_{f,h}(r)$ is defined as follows.
Let $W(\cdot)$ denote a standard Brownian motion on $[0,1]$. For $h\in(0,\infty)$, define $f_{h}(r)=(1,\exp(-hr))^{\prime}$, and for $h=0$, let $f_{h}(r)=(1,r)^{\prime}$. Define
$I_{h}(r)$ $:=\begin{cases}\int_{0}^{r}\exp(-(r-s)h)dW(s)&\text{for }h>0\\ W(r)&\text{for }h=0\end{cases}\quad\text{ and}$
$I_{f,h}(r)$ $:=I_{h}(r)-\left(\int_{0}^{1}I_{h}(s)f_{h}(s)ds\right)^{\prime}\left(\int_{0}^{1}f_{h}(s)f_{h}(s)^{\prime}ds\right)^{-1}f_{h}(r),$ (10)
where the subscript $f$ indicates that we take the residual process of $I_{h}(\cdot)$ after projecting it onto the space spanned by $f_{h}(\cdot)$. When $h>0$,
$I_{f,h}(r)=\$$I_{h}(r)-\alpha_{h}(r)\int_{0}^{1}I_{h}(s)ds-\beta_{h}(r)\int_{0}^{1}\frac{1-e^{-hs}}{h}I_{h}(s)ds,$ (11)
where
$\alpha_{h}(r)$ $=\ \frac{h[1-e^{-2h}-2(1-e^{-h})e^{-hr}+2he^{-hr}-2(1-e^{-h})]}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\quad\text{and}$
$\beta_{h}(r)$ $=\ \frac{2h^{2}[he^{-hr}-(1-e^{-h})]}{h(1-e^{-2h})-2(1-e^{-h})^{2}}.$ (12)
When $h=0$, $I_{f,h}(r)$ reduces to detrended Brownian motion:
$W_{f}(r):=W(r)-(4-6r)\int_{0}^{1}W(s)ds-(12r-6)\int_{0}^{1}sW(s)ds.$ (13)
The following lemma establishes the connection between $I_{f,h}(r)$ for $h>0$ and $h=0$.
Lemma 1.
Let $h\downarrow 0$, we have
$\sup_{r\in[0,1]}|\alpha_{h}(r)-(4-6r)|\to 0\quad\text{and}\quad\sup_{r\in[0,1]}|\beta_{h}(r)-(12r-6)|\to 0.$
Remark 2.
This lemma shows that $I_{f,h}(r)$ is uniformly continuous with respect to $h$. Consequently, $I_{f,h}(r)\to I_{f,0}(r)$ uniformly over $r\in[0,1]$ as $h\to 0$ a.s. and the asymptotic distribution is continuous at $h=0$.
Supplemental Material
to
Initial-Condition-Robust Inference
in Autoregressive Models
Donald W. K. Andrews***Andrews: Department of Economics and Cowles Foundation, Yale University. donald.andrews@yale.edu.    Ming LiLi: Department of Economics and Risk Management Institute, National University of Singapore. mli@nus.edu.sg.    Yapeng ZhengZheng: Department of Economics, Chinese University of Hong Kong. yapengzheng@link.cuhk.edu.hk.
February 7, 2026
In this Supplemental Material, Section A describes how the critical values are computed, Section B proves Theorem 1, Section C proves Theorem 2, and Section D provides some additional simulation results.
Appendix A Computation of Critical Values
For a given $\alpha$, we compute $c_{h}(\alpha)$ by simulating the asymptotic distribution of $J_{h}$. To do so, we simulate $B=300,000$ sample paths of a standard Brownian motion on the interval $[0,1]$ by discretizing the interval into $N=50,000$ equal subintervals with time step $1/N$. We initialize the process at $W(0)=0$ and construct each path iteratively as
$W(r_{i})=W(r_{i-1})+Z_{i-1}/N^{1/2},\ \ r_{i}=i/N,\ i=1,...,N,$ (14)
where $Z_{i-1}$ are independent standard normal random variables. Next, we compute $I_{h}(r_{j})$, $j=1,\dots,N$, using (10) via a numerical approximation:
$I_{h}(r_{j})=\begin{cases}\sum_{i=1}^{j}\exp(-(r_{j}-r_{i})h)\cdot(W(r_{i})-W(r_{i-1}))&\text{for }h>0,\\ W(r_{j})&\text{for }h=0.\end{cases}$ (15)
Then, we use (10) to compute $I_{f,h}(r_{j})$, $j=1,\dots,N$. Given $I_{f,h}(r_{j})$ and $W(r_{j})$, we evaluate $J_{h}$ in (9) via numerical approximation. Finally, the simulated estimate of $c_{h}(\alpha)$ is obtained as the $\alpha$-th quantile of the empirical distribution of the $B$ realizations of $J_{h}$.
Appendix B Proof of Theorem 1
To prove Theorem 1, it is sufficient to verify Assumptions B1* and B2 of [2] for $CS_{n}=CI_{\text{ICR},n}$. In the present case $\lambda=(\rho,F)$, $r(\lambda)=\rho$, $h_{n}(\lambda)=n(1-\rho)\in R$, $H=[0,\infty]$, and the parameter space $\Lambda$ is defined in Section 5.1.
For Assumption B1*, consider a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda:n\geq 1\}$ for which $h_{n}(\lambda_{n})\to h\in H$, i.e., $\rho_{n}=1-h_{n}/n$ and $h_{n}\to h\in[0,\infty]$. We have
$CP_{n}(\lambda_{n})=P_{\lambda_{n}}(\rho_{n}\in CI_{\text{ICR},n})=P_{\lambda_{n}}(c_{h_{n}}(\alpha/2)\leq T_{n}(\rho_{n})\leq c_{h_{n}}(1-\alpha/2)).$ (16)
By Theorem 2, we have $T_{n}(\rho_{n})\to_{d}J_{h}$ under $\{\lambda_{n}\in\Lambda_{n}:n\geq 1\}$. In addition, $c_{h_{n}}(\beta)\to c_{h}(\beta)$ for $\beta=\alpha/2$ and $1-\alpha/2$. This is true because the distribution function of $J_{h}$ is strictly increasing at its $\beta$-th quantile. Thus, for any $\varepsilon>0$, we have
$L_{n}(c_{h}(\beta)-\varepsilon)$ $\to F_{J_{h}}(c_{h}(\beta)-\varepsilon)<\beta\text{ and }L_{n}(c_{h}(\beta)+\varepsilon)\to F_{J_{h}}(c_{h}(\beta)+\varepsilon)>\beta,$ (17)
where $L_{n}(x)$ denotes the distribution function of $T_{n}(\rho_{n})$ at $x$ and $F_{J_{h}}(x)$ denotes the distribution function of $J_{h}$ at $x$. This and the definition $c_{h_{n}}(\beta)=\inf\{x\in R:F_{J_{h_{n}}}(x)\geq\beta\}$ yield $1\{c_{h}(\beta)-\varepsilon\leq c_{h_{n}}(\beta)\leq c_{h}(\beta)+\varepsilon\}\to 1$ as $n\to\infty$ for any $\varepsilon>0$. By the definition of convergence in distribution and continuity of $J_{h}$, it follows that $CP_{n}(\lambda_{n})\to 1-\alpha$. Assumption B1* therefore holds with $CP=1-\alpha$ for all $h\in H$.
For Assumption B2, assume we are given $\{\lambda_{p_{n}}\in\Lambda:n\geq 1\}$ for a subsequence $\{p_{n}\}_{n\geq 1}$ of $\{n\}_{n\geq 1}$ such that $h_{p_{n}}(\lambda_{p_{n}})\to h\in H$. Define $\{\lambda_{n}^{*}:n\geq 1\}$ by (i) $\lambda_{p_{n}}^{*}=\lambda_{p_{n}}$ $\forall n\geq 1$, (ii) when $h<\infty$ and $m\neq p_{n}$, define $\lambda_{m}^{*}=(1-h/m,F^{*})$, and (iii) when $h=\infty$ and $m\neq p_{n}$, define $\lambda_{m}^{*}=(0,F^{*})$, where $F^{*}$ is the distribution such that $\{U_{i}:i=1,2,\dots\}$ are i.i.d. standard normal. Then, $\lambda_{n}^{*}\in\Lambda$ for all $n\geq 1$ and by construction $h_{n}(\lambda_{n}^{*})\to h\in H$. This verifies Assumption B2 and completes the proof.
Appendix C Proof of Theorem 2
C.1.  Technical Lemmas
Let
$\widetilde{U}$ $=(U_{1},U_{2}+\rho_{n}U_{1},\dots,\sum_{j=1}^{n}\rho_{n}^{n-j}U_{j})^{\prime},$
$\widetilde{U}_{-1}$ $=(0,U_{1},\dots,\sum_{j=1}^{n-1}\rho_{n}^{n-1-j}U_{j})^{\prime},\text{ and }\widetilde{U}_{0}=\sum_{j=0}^{\infty}\rho_{n}^{j}U_{-j}.$ (18)
When $\rho_{n}\neq 1$, define $S=(1-\rho_{n}^{n})/(1-\rho_{n})$ and $Q=(1-\rho_{n}^{2n})/(1-\rho_{n}^{2})$. Let $[x]$ denote the integer part of $x$.
The next lemma provides several useful results for the regime $h_{n}\coloneqq n(1-\rho_{n})\to h\in(0,\infty)$. Unless stated otherwise, we let $n\rightarrow\infty$ in presenting the asymptotic results.
Lemma 2.
When $n(1-\rho_{n})\to h\in(0,\infty)$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
  1. (a)
    $Q=\frac{1-e^{-2h}}{2h}n+o(n)$,
  2. (b)
    $S=\frac{1-e^{-h_{n}}}{h_{n}}n+o(n)$,
  3. (c)
    $\frac{n^{2}}{nQ-S^{2}}\to\frac{2h^{2}}{h(1-e^{-2h})-2(1-e^{-h})^{2}}$ ,
  4. (d)
    $\frac{nS}{nQ-S^{2}}\to\frac{2h(1-e^{-h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}$,
  5. (e)
    $\frac{nQ}{nQ-S^{2}}\to\frac{h(1-e^{-2h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}$,
  6. (f)
    $n^{-1}\sum_{i=1}^{n}\left(\frac{n(Q-S\rho^{i-1})}{nQ-S^{2}}\right)^{2}=\frac{nQ}{nQ-S^{2}}$,
  7. (g)
    $n^{-1}\sum_{i=1}^{n}\frac{n^{2}(Q-S\rho^{i-1})(n\rho^{i-1}-S)}{(nQ-S^{2})^{2}}=\frac{-nS}{(nQ-S^{2})}$, and
  8. (h)
    $n^{-1}\sum_{i=1}^{n}\left(\frac{n(n\rho^{i-1}-S)}{nQ-S^{2}}\right)^{2}=\frac{n^{2}}{nQ-S^{2}}$.
The next lemma provides several useful results for the regime $n(1-\rho_{n})\to 0$.
Lemma 3.
When $n(1-\rho_{n})\to h=0$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
  1. (a)
    $\frac{1-\rho_{n}^{i-1}}{n(1-\rho_{n})}\to r$ if $i=[nr]$,
  2. (b)
    $\frac{n(Q-2S+n)}{nQ-S^{2}}\to 4$,
  3. (c)
    $\frac{n^{2}(1-\rho_{n})(n-S)}{nQ-S^{2}}\to 6$, and
  4. (d)
    $\frac{n^{4}(1-\rho_{n})^{2}}{nQ-S^{2}}\to 12$.
The next lemma provides a set of convergence results that are used in the proof of Theorem 2.
Lemma 4.
When $n(1-\rho_{n})\to h\in(0,\infty)$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
  1. (a)
    $n^{-1/2}\widetilde{U}_{[nr]}\Rightarrow I_{h}(r)$,
  2. (b)
    $n^{-1}\sum_{i=1}^{n}U_{i}\stackrel{{\scriptstyle p}}{{\to}}0$,
  3. (c)
    $n^{-1}\sum_{i=1}^{n}U_{i}^{2}\stackrel{{\scriptstyle p}}{{\to}}1$,
  4. (d)
    $n^{-1/2}\sum_{i=1}^{n}U_{i}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}dW(r)$,
  5. (e)
    $n^{-1/2}\sum_{i=1}^{n}\rho_{n}^{i-1}U_{i}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}e^{-hr}dW(r)$,
  6. (f)
    $n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}I_{h}(r)dr$,
  7. (g)
    $n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}e^{-hr}I_{h}(r)dr$,
  8. (h)
    $n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}I_{h}(r)dW(r)$,
  9. (i)
    $n^{-1}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}U_{i}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}e^{-hr}I_{h}(r)dW(r)$,
  10. (j)
    $n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}I_{h}^{2}(r)dr$,
  11. (k)
    $n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}^{2}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}I_{h}(r)dr$,
  12. (l)
    $n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}U_{i}^{2}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}e^{-hr}I_{h}(r)dr$,
  13. (m)
    $n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}U_{i}^{2}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}I_{h}^{2}(r)dr$, and
  14. (n)
    $n^{-1-l_{1}/2}\sum_{i=2}^{n}\left(\rho_{n}^{i-1}\right)^{l_{0}}\widetilde{U}_{i-1}^{l_{1}}U_{i}^{l_{2}}=o_{p}(n)$ for $(l_{0},l_{1},l_{2})=(0,1,0)$, $(0,1,1)$, $(0,2,0)$, $(0,2,1)$, $(0,3,0)$, $(0,3,1)$, $(0,4,0)$, $(1,1,1)$, $(1,2,0)$, $(1,3,0)$, $(2,2,0)$, and $(1,3,0)$.
The following three lemmas are needed for the proof of the case $h=\infty$ in Theorem 2. Since they are simplified versions of Lemmas 6, 7, and 9 in [3, AG12 hereafter], respectively, we omit their proofs.
Lemma 5.
When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
$E(\widetilde{U}_{0}^{2}U_{1}^{2})-(1-\rho_{n}^{2})^{-1}(EU_{1}^{2})^{2}$ $=O(1),$
$E\widetilde{U}_{0}^{2}-(1-\rho_{n}^{2})^{-1}EU_{1}^{2}$ $=O(1),$
$E\widetilde{U}_{0}$ $=O(1),\quad\text{and}$
$E(\widetilde{U}_{0}U_{1}^{2})$ $=O(1).$
Lemma 6.
When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, we have
$E\left(\sum_{i=2}^{n}\left[E\zeta_{i}^{2}-E(\zeta_{i}^{2}|\mathcal{G}_{i-1})\right]\right)^{2}\to 0,\quad\text{where }\zeta_{i}\equiv n^{-1/2}\frac{\widetilde{U}_{i-1}U_{i}}{(E(\widetilde{U}_{0}^{2}U_{1}^{2}))^{1/2}}.$
Lemma 7.
When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, we have
$\sum_{i=2}^{n}E(\zeta_{i}^{2}1(|\zeta_{i}|>\delta)|\mathcal{G}_{i-1})\stackrel{{\scriptstyle p}}{{\to}}0$
for any $\delta>0$.
The next lemma is slightly different from Lemma 8 in AG12, so we prove the new results. In this lemma, we redefine $X_{2}(\rho_{n})$ as an $n\times 2$ matrix whose $i$-th row is $(1,n^{1/2}(1-\rho_{n})^{1/2}\rho_{n}^{i-1})$ and let $X(\rho_{n})=[X_{1}:X_{2}(\rho_{n})]$. Note that this redefinition does not affect $P_{X_{2}(\rho_{n})}$ or $M_{X_{2}(\rho_{n})}$.
Lemma 8.
When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly, we have
  1. (a)
    $n^{-1}(1-\rho_{n})^{1/2}\widetilde{U}_{-1}^{\prime}X_{2}(\rho_{n})=(o_{p}(1),o_{p}(1))$,
  2. (b)
    $(E\widetilde{U}_{0}^{2})^{-1}n^{-1}\widetilde{U}_{-1}^{\prime}\widetilde{U}_{-1}\stackrel{{\scriptstyle p}}{{\to}}1$,
  3. (c)
    $(E(\widetilde{U}_{0}^{2}U_{1}^{2}))^{-1}n^{-1}\sum_{i=1}^{n}\widetilde{U}_{i-1}^{2}U_{i}^{2}\stackrel{{\scriptstyle p}}{{\to}}1$,
  4. (d)
    $(X(\rho_{n})^{\prime}X(\rho_{n}))^{-1}X(\rho_{n})^{\prime}U=(O_{p}(n^{-1/2}(1-\rho_{n})^{3/2}),O_{p}(n^{-1/2}(1-\rho_{n})),O_{p}(n^{-1/2}(1-\rho_{n})))^{\prime}$,
  5. (e)
    $(E(\widetilde{U}_{0}^{2}U_{1}^{2}))^{-1}n^{-1}\widetilde{U}_{-1}^{\prime}\Delta^{2}\widetilde{U}_{-1}\stackrel{{\scriptstyle p}}{{\to}}1$,
  6. (f)
    $(1-\rho_{n})^{1/2}n^{-1}(X_{2}(\rho_{n})^{\prime}\Delta^{2}\widetilde{U}_{-1})=O_{p}(1)$, and
  7. (g)
    $n^{-1}(X_{2}(\rho_{n})^{\prime}\Delta^{2}X_{2}(\rho_{n}))=O_{p}(1)$.
C.2.  Proof of Theorem 2
C.2.1 Local-to-Unity Case
We first handle the case for $h\in(0,\infty)$. We can write
$n(\widehat{\rho}_{n}(\rho_{n})-\rho_{n})=\$$(n^{-2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}X_{1})^{-1}n^{-1}X_{1}^{\prime}M_{X_{2}(\rho_{n})}U\quad\text{and}$
$n\widehat{\sigma}_{n}^{2}(\rho_{n})=\$$(n^{-2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}X_{1})^{-1}(n^{-2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}\Delta^{2}M_{X_{2}(\rho_{n})}X_{1})$
$\times(n^{-2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}X_{1})^{-1}.$ (19)
We consider the terms in (19) one at a time. First, we have
$M_{X_{2}(\rho_{n})}X_{1}=M_{X_{2}(\rho_{n})}(X_{2}(\rho_{n})(\mu,Y_{0}^{*})^{\prime}+\widetilde{U}_{-1})=M_{X_{2}(\rho_{n})}\widetilde{U}_{-1}.$ (20)
In consequence,
$n^{-2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}X_{1}$
[…]
$+\frac{2nS}{nQ-S^{2}}\left(n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}\right)\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)$
$-\frac{n^{2}}{nQ-S^{2}}\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)^{2}$
$\stackrel{{\scriptstyle d}}{{\to}}\$$\int_{0}^{1}I_{h}^{2}(r)dr-\frac{h(1-e^{-2h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}I_{h}(r)dr\right)^{2}$
$+\frac{4h(1-e^{-h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}I_{h}(r)dr\right)\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)$
$-\frac{2h^{2}}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)^{2}$
$=\$$\int_{0}^{1}I_{f,h}^{2}(r)dr.$ (21)
where the convergence holds by Lemma 2 and parts (f) and (g) of Lemma 4 and the last equality holds by (10).
Similarly, we have
$n^{-1}X_{1}^{\prime}M_{X_{2}(\rho_{n})}U$
[…]
$+\frac{nS}{nQ-S^{2}}\left(n^{-1/2}\sum_{i=1}^{n}U_{i}\right)\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)$
$+\frac{nS}{nQ-S^{2}}\left(n^{-1/2}\sum_{i=1}^{n}\rho_{n}^{i-1}U_{i}\right)\left(n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}\right)$
$-\frac{n^{2}}{nQ-S^{2}}\left(n^{-1/2}\sum_{i=1}^{n}\rho_{n}^{i-1}U_{i}\right)\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)$
$\stackrel{{\scriptstyle d}}{{\to}}\$$\int_{0}^{1}I_{h}(r)dW(r)-\frac{h(1-e^{-2h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}dW(r)\right)\left(\int_{0}^{1}I_{h}(r)dr\right)$
$+\frac{2h(1-e^{-h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}dW(r)\right)\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)$
$+\frac{2h(1-e^{-h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}e^{-hr}dW(r)\right)\left(\int_{0}^{1}I_{h}(r)dr\right)$
$-\frac{2h^{2}}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}e^{-hr}dW(r)\right)\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)$
$=\$$\int_{0}^{1}I_{f,h}(r)dW(r),$ (22)
where the convergence uses Lemma 2 and parts (d)(g) of Lemma 4.
To determine the asymptotic distribution of $n^{-2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}\Delta^{2}M_{X_{2}(\rho_{n})}X_{1}$, we proceed as follows. The space spanned by $X(\rho_{n})$ is equivalent to the space spanned by $\{(1,\rho_{n}^{i-1},n^{-1/2}\widetilde{U}_{i-1})\}_{i=1}^{n}$. Therefore, we have
$\widehat{U}_{i}$ $=U_{i}-A_{n}^{\prime}B_{n}^{-1}C_{n,i},\text{ where}$
$A_{n}$ $=\begin{pmatrix}n^{-1/2}\sum_{j=1}^{n}U_{j}\\ n^{-1/2}\sum_{j=1}^{n}\rho_{n}^{j-1}U_{j}\\ n^{-1}\sum_{j=2}^{n}\widetilde{U}_{j-1}U_{j}\end{pmatrix}$,
$B_{n}$ $=\begin{pmatrix}1&n^{-1}S&n^{-3/2}\sum_{j=2}^{n}\widetilde{U}_{j-1}\\ n^{-1}S&n^{-1}Q&n^{-3/2}\sum_{j=2}^{n}\rho_{n}^{j-1}\widetilde{U}_{j-1}\\ n^{-3/2}\sum_{j=2}^{n}\widetilde{U}_{j-1}&n^{-3/2}\sum_{j=2}^{n}\rho_{n}^{j-1}\widetilde{U}_{j-1}&n^{-2}\sum_{j=2}^{n}\widetilde{U}_{j-1}^{2}\end{pmatrix},\text{ and}$
$C_{n,i}$ $=\begin{pmatrix}n^{-1/2}\\ n^{-1/2}\rho_{n}^{i-1}\\ n^{-1}\widetilde{U}_{i-1}\end{pmatrix}$. (23)
For each term in $A_{n}$, we have
$n^{-1/2}\sum_{j=1}^{n}U_{j}$ $\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}dW(r),$
$n^{-1/2}\sum_{j=1}^{n}\rho_{n}^{j-1}U_{j}$ $\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}e^{-hr}dW(r),\text{ and}$
$n^{-1}\sum_{j=2}^{n}\widetilde{U}_{j-1}U_{j}$ $=\int_{0}^{1}I_{h}(r)dW(r)$ (24)
by Lemma 4(d), (e), and (h), respectively. Therefore, $A_{n}=O_{p}(1)$.
For the entries of $B_{n}$, first we have
$n^{-1}S$ $=\frac{1-e^{-h}}{h}+o(1)=O(1)\text{ and}$
$n^{-1}Q$ $=\frac{1-e^{-2h}}{2h}+o(1)=O(1)$ (25)
by Lemma 2(a) and (b). Using Lemma 4(f), (g), and (j), we have
$n^{-3/2}\sum_{j=2}^{n}\widetilde{U}_{j-1}$ $\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}I_{h}(r)dr,$
$n^{-3/2}\sum_{j=2}^{n}\rho_{n}^{j-1}\widetilde{U}_{j-1}$ $\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}e^{-hr}dW(r),\text{ and}$
$n^{-2}\sum_{j=2}^{n}\widetilde{U}_{j-1}^{2}$ $\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}I_{h}^{2}(r)dr.$ (26)
Therefore, $B_{n}=O_{p}(1)$. Also, note that
$\frac{n^{2}}{nQ-S^{2}}\det(B_{n})=\$$n^{-2}\sum_{j=2}^{n}\widetilde{U}_{j-1}^{2}-\frac{n^{2}}{nQ-S^{2}}\left(n^{-3/2}\sum_{j=2}^{n}\rho_{n}^{j-1}\widetilde{U}_{j-1}\right)^{2}$
$+2\frac{nS}{nQ-S^{2}}\left(n^{-3/2}\sum_{j=2}^{n}\widetilde{U}_{j-1}\right)\left(n^{-3/2}\sum_{j=2}^{n}\rho_{n}^{j-1}\widetilde{U}_{j-1}\right)$
$-\frac{nQ}{nQ-S^{2}}\left(n^{-3/2}\sum_{j=2}^{n}\widetilde{U}_{j-1}\right)^{2}$
$\stackrel{{\scriptstyle d}}{{\to}}\$$\int_{0}^{1}I_{h}^{2}(r)dr-\frac{2h^{2}}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)^{2}$
$+\frac{4h(1-e^{-h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}I_{h}(r)dr\right)\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)$
$-\frac{h(1-e^{-2h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}I_{h}(r)dr\right)^{2}$
$=\$$\int_{0}^{1}I_{f,h}^{2}(r)dr.$ (27)
Since $P(\int_{0}^{1}I_{f,h}^{2}(r)dr=0)=0$, we have $P(\det(B_{n})=0)\to 0$ as $n\to\infty$, which implies that $B_{n}$ is invertible with probability approaching one. Combining it with the fact that $B_{n}=O_{p}(1)$, we have $B_{n}^{-1}=O_{p}(1)$.
Next, we define
$D_{n}$ $=\begin{pmatrix}n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}U_{i}\\ n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}\rho_{n}^{i-1}U_{i}\\ n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}U_{i}\end{pmatrix}\text{ and}$
$K_{n}$ $=\begin{pmatrix}n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}&n^{-2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}&n^{-5/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}\\ n^{-2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}&n^{-2}\sum_{i=2}^{n}\rho_{n}^{2i-2}\widetilde{U}_{i-1}^{2}&n^{-5/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}\\ n^{-5/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}&n^{-5/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}&n^{-3}\sum_{i=2}^{n}\widetilde{U}_{j-1}^{4}\end{pmatrix}$. (28)
Using the results of Lemma 4(n) with $(l_{0},l_{1},l_{2})=(0,2,1)$, $(1,1,1)$, and $(0,3,1)$, we have $D_{n}=o_{p}(n)$. Similarly, using Lemma 4(j) and Lemma 4(n) with $(l_{0},l_{1},l_{2})=(2,2,0)$, $(0,3,0)$, $(1,3,0)$, and $(0,4,0)$, we have $K_{n}=o_{p}(n)$.
By (23), we have
$n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}\widehat{U}_{i}^{2}$ $=n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}U_{i}^{2}-2n^{-1}A_{n}^{\prime}B_{n}^{-1}D_{n}+n^{-1}A_{n}^{\prime}B_{n}^{-1}K_{n}B_{n}^{-1}A_{n}$
$=n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}U_{i}^{2}+o_{p}(1),$ (29)
where the second equality holds by $A_{n}=O_{p}(1)$, $B_{n}^{-1}=O_{p}(1)$, $D_{n}=o_{p}(n)$, and $K_{n}=o_{p}(n)$.
Proceeding analogously to (29), replacing $\widetilde{U}_{i-1}^{2}$ with $\widetilde{U}_{i-1}$ and then omitting the $\widetilde{U}_{i-1}^{2}$ term, we have
$n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}\widehat{U}_{i}^{2}$ $=n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}^{2}+o_{p}(1)\quad\text{and}$
$n^{-1}\sum_{i=1}^{n}c_{i}\widehat{U}_{i}^{2}$ $=n^{-1}\sum_{i=1}^{n}c_{i}U_{i}^{2}+o_{p}(1)$ (30)
for any non-random sequence $c_{i}=O(1)$.
Next, we have
$n^{-2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}\Delta^{2}M_{X_{2}(\rho_{n})}X_{1}$
[…]
$+n^{-2}\frac{1}{(nQ-S^{2})^{2}}\sum_{i=1}^{n}\left(\sum_{k=2}^{n}\left(Q-S(\rho_{n}^{k-1}+\rho_{n}^{i-1})+n\rho^{i+k-2}\right)\widetilde{U}_{k-1}\right)^{2}\widehat{U}_{i}^{2}+o_{p}(1)$
[…]
where the second equality holds by $p_{ii}^{*}=\min\{p_{ii},n^{-1/2}\}=O_{p}(n^{-1/2})$. We now analyze $\mathcal{U}_{1,n}$, $\mathcal{U}_{2,n}$, and $\mathcal{U}_{3,n}$ in turn.
For $\mathcal{U}_{1,n}$, by (29) and Lemma 4(m), we have
$\mathcal{U}_{1,n}=n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}U_{i}^{2}+o_{p}(1)\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}I_{h}^{2}(r)dr.$ (32)
For $\mathcal{U}_{2,n}$, we have
$\mathcal{U}_{2,n}=\$$\frac{nQ}{nQ-S^{2}}\left(n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}\right)\left(n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}\widehat{U}_{i}^{2}\right)$
$-\frac{nS}{nQ-S^{2}}\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)\left(n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}\widehat{U}_{i}^{2}\right)$
$-\frac{nS}{nQ-S^{2}}\left(n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}\right)\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\widehat{U}_{i}^{2}\right)$
$+\frac{n^{2}}{nQ-S^{2}}\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\widehat{U}_{i}^{2}\right)$
[…]
$-\frac{nS}{nQ-S^{2}}\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)\left(n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}^{2}\right)$
$-\frac{nS}{nQ-S^{2}}\left(n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}\right)\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}U_{i}^{2}\right)$
$+\frac{n^{2}}{nQ-S^{2}}\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)\left(n^{-3/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}U_{i}^{2}\right)+o_{p}(1)$
$\stackrel{{\scriptstyle d}}{{\to}}\$$\frac{h(1-e^{-2h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}I_{h}(r)dr\right)^{2}$
$-\frac{4h(1-e^{-h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)\left(\int_{0}^{1}I_{h}(r)dr\right)$
$+\frac{2h^{2}}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)^{2},$ (33)
where the first equality holds by the definition of $\mathcal{U}_{2,n}$ and algebra, the second equality holds by (30), and the convergence holds by Lemma 2(c)(e) and Lemma 4(f), (g), (k), and (l).
For $\mathcal{U}_{3,n}$, first notice that
$\left(\sum_{k=2}^{n}\left(Q-S(\rho_{n}^{k-1}+\rho_{n}^{i-1})+n\rho^{i+k-2}\right)\widetilde{U}_{k-1}\right)^{2}$
$=\$$\left((Q-S\rho^{i-1})\left(\sum_{k=2}^{n}\widetilde{U}_{k-1}\right)+(n\rho^{i-1}-S)\left(\sum_{k=2}^{n}\rho_{n}^{k-1}\widetilde{U}_{k-1}\right)\right)^{2}$
$=\$$(Q-S\rho^{i-1})^{2}\left(\sum_{k=2}^{n}\widetilde{U}_{k-1}\right)^{2}+2(Q-S\rho^{i-1})(n\rho^{i-1}-S)\left(\sum_{k=2}^{n}\widetilde{U}_{k-1}\right)\left(\sum_{k=2}^{n}\rho_{n}^{k-1}\widetilde{U}_{k-1}\right)$
$+(n\rho^{i-1}-S)^{2}\left(\sum_{k=2}^{n}\rho_{n}^{k-1}\widetilde{U}_{k-1}\right)^{2}.$ (34)
Thus, we have
$\mathcal{U}_{3,n}=\$$\left(n^{-3/2}\sum_{k=2}^{n}\widetilde{U}_{k-1}\right)^{2}\left(n^{-1}\sum_{i=1}^{n}\left(\frac{n(Q-S\rho^{i-1})}{nQ-S^{2}}\right)^{2}\widehat{U}_{i}^{2}\right)$
$+\left(n^{-3/2}\sum_{k=2}^{n}\widetilde{U}_{k-1}\right)\left(n^{-3/2}\sum_{k=2}^{n}\rho_{n}^{k-1}\widetilde{U}_{k-1}\right)\left(2n^{-1}\sum_{i=1}^{n}\frac{n^{2}(Q-S\rho^{i-1})(n\rho^{i-1}-S)}{(nQ-S^{2})^{2}}\widehat{U}_{i}^{2}\right)$
$+\left(n^{-3/2}\sum_{k=2}^{n}\rho_{n}^{k-1}\widetilde{U}_{k-1}\right)^{2}\left(n^{-1}\sum_{i=1}^{n}\left(\frac{n(n\rho^{i-1}-S)}{nQ-S^{2}}\right)^{2}\widehat{U}_{i}^{2}\right)$
$=\$$\left(n^{-3/2}\sum_{k=2}^{n}\widetilde{U}_{k-1}\right)^{2}\left(n^{-1}\sum_{i=1}^{n}\left(\frac{n(Q-S\rho^{i-1})}{nQ-S^{2}}\right)^{2}U_{i}^{2}\right)$
$+\left(n^{-3/2}\sum_{k=2}^{n}\widetilde{U}_{k-1}\right)\left(n^{-3/2}\sum_{k=2}^{n}\rho_{n}^{k-1}\widetilde{U}_{k-1}\right)\left(2n^{-1}\sum_{i=1}^{n}\frac{n^{2}(Q-S\rho^{i-1})(n\rho^{i-1}-S)}{(nQ-S^{2})^{2}}U_{i}^{2}\right)$
$+\left(n^{-3/2}\sum_{k=2}^{n}\rho_{n}^{k-1}\widetilde{U}_{k-1}\right)^{2}\left(n^{-1}\sum_{i=1}^{n}\left(\frac{n(n\rho^{i-1}-S)}{nQ-S^{2}}\right)^{2}U_{i}^{2}\right)+o_{p}(1)$
$\stackrel{{\scriptstyle d}}{{\to}}\$$\frac{h(1-e^{-2h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}I_{h}(r)dr\right)^{2}$
$-\frac{4h(1-e^{-h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}I_{h}(r)dr\right)\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)$
$+\frac{2h^{2}}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)^{2},$ (35)
where the second equality holds by (30) and the convergence holds by Lemma 2(f)(h) and Lemma 4(c), (f), and (g).
Combining (32), (33), and (35), we have
$\mathcal{U}_{1,n}-2\mathcal{U}_{2,n}+\mathcal{U}_{3n}$
$\stackrel{{\scriptstyle d}}{{\to}}\$$\int_{0}^{1}I_{h}(r)^{2}dr-\frac{h(1-e^{-2h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}I_{h}(r)dr\right)^{2}$
$+\frac{4h(1-e^{-h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}I_{h}(r)dr\right)\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)$
$-\frac{2h^{2}}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)^{2}$
$=\$$\int_{0}^{1}I_{f,h}^{2}(r)dr.$ (36)
Thus, we have
$n^{-2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}\Delta^{2}M_{X_{2}(\rho_{n})}X_{1}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}I_{f,h}^{2}(r)dr.$ (37)
Putting (21), (22), and (37) together gives
$T_{n}(\rho_{n})\stackrel{{\scriptstyle d}}{{\to}}\frac{\int_{0}^{1}I_{f,h}(r)dW(r)}{\left(\int_{0}^{1}I_{f,h}(r)^{2}dr\right)^{1/2}}.$ (38)
C.2.2 Close-to-Unit-Root/Unit-Root Case
Next, we consider the case $h=0$. We rewrite (21) as
$n^{-2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}X_{1}$
[…]
$\times\sum_{k=2}^{n}\Bigg(Q-2S+n-n(1-\rho_{n})(n-S)\left(\frac{1-\rho_{n}^{i-1}}{n(1-\rho_{n})}+\frac{1-\rho_{n}^{k-1}}{n(1-\rho_{n})}\right)$
$+n^{3}(1-\rho_{n})^{2}\frac{1-\rho_{n}^{i-1}}{n(1-\rho_{n})}\frac{1-\rho_{n}^{k-1}}{n(1-\rho_{n})}\Bigg)\widetilde{U}_{k-1}\Bigg)$
[…]
$\times\sum_{k=2}^{n}\Bigg(Q-2S+n-n(1-\rho_{n})(n-S)\left(\frac{1-\rho_{n}^{i-1}}{n(1-\rho_{n})}+\frac{1-\rho_{n}^{k-1}}{n(1-\rho_{n})}\right)$
$+n^{3}(1-\rho_{n})^{2}\frac{1-\rho_{n}^{i-1}}{n(1-\rho_{n})}\frac{1-\rho_{n}^{k-1}}{n(1-\rho_{n})}\Bigg)\widetilde{U}_{k-1}$
[…]
$+\frac{2n^{2}(1-\rho_{n})(n-S)}{nQ-S^{2}}\left(n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}\right)\left(n^{-3/2}\sum_{i=2}^{n}\frac{1-\rho_{n}^{i-1}}{n(1-\rho_{n})}\widetilde{U}_{i-1}\right)$
$-\frac{n^{4}(1-\rho_{n})^{2}}{nQ-S^{2}}\left(n^{-3/2}\sum_{i=2}^{n}\frac{1-\rho_{n}^{i-1}}{n(1-\rho_{n})}\widetilde{U}_{i-1}\right)^{2}$
$\stackrel{{\scriptstyle d}}{{\to}}\$$\int_{0}^{1}W^{2}(r)dr-4\left(\int_{0}^{1}W(r)dr\right)^{2}+12\left(\int_{0}^{1}W(r)dr\right)\left(\int_{0}^{1}rW(r)dr\right)$
$-12\left(\int_{0}^{1}rW(r)dr\right)^{2}$
$=\$$\int_{0}^{1}W_{f}^{2}(r)dr,$ (39)
where the convergence holds by Lemma 3(a)(d). Observe that
$n^{-1}X_{1}^{\prime}M_{X_{2}(\rho_{n})}U\ \text{and}\ n^{-2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}\Delta^{2}M_{X_{2}(\rho_{n})}X_{1}$ (40)
can be treated analogously by rewriting (22) and (31) in the form of (39). Therefore, we have
$n^{-1}X_{1}^{\prime}M_{X_{2}(\rho_{n})}U$ $\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}W_{f}(r)dW(r)\text{ and }$
$n^{-2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}\Delta^{2}M_{X_{2}(\rho_{n})}X_{1}$ $\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}W_{f}^{2}(r)dr.$ (41)
Combining (39)–(41), we have
$T_{n}(\rho_{n})\stackrel{{\scriptstyle d}}{{\to}}\frac{\int_{0}^{1}W_{f}(r)dW(r)}{\left(\int_{0}^{1}W_{f}^{2}(r)dr\right)^{1/2}}$ (42)
when $h=0$.
Remark 3.
Our proof for $h=0$ differs from that of Theorem 1 in AG12, where it is necessary to show that the influence of $Y_{0}^{*}$ is negligible. In contrast, we eliminate the influence of $Y_{0}^{*}$ via the projection approach and thus avoid imposing any conditions on $Y_{0}^{*}$.
C.2.3 Stationary Case
It remains to analyze the case $h=\infty$. It suffices to consider the two sub-cases $\rho_{n}\to\rho^{*}<1$ and $\rho_{n}\to 1$. First, assume that $\rho_{n}\to 1$ and $n(1-\rho_{n})\to\infty$. We define
$a_{n}=n^{1/2}\frac{E\widetilde{U}_{0}^{2}}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}\quad\text{and}\quad d_{n}=\frac{E\widetilde{U}_{0}^{2}}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}.$ (43)
We redefine $X_{2}(\rho_{n})$ as an $n\times 2$ matrix whose $i$-th row is $(1,h_{n}^{1/2}\rho_{n}^{i-1})$. This redefinition does not affect $P_{X_{2}(\rho_{n})}$ or $M_{X_{2}(\rho_{n})}$, but simplifies the analysis. By algebra, we have
$a_{n}(\widehat{\rho}_{n}(\rho_{n})-\rho_{n})$ $=\left(n^{-1}\frac{X_{1}^{\prime}M_{X_{2}(\rho_{n})}X_{1}}{E\widetilde{U}_{0}^{2}}\right)^{-1}\frac{n^{-1/2}X_{1}^{\prime}M_{X_{2}(\rho_{n})}U}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}$
$=\left(n^{-1}\frac{\widetilde{U}_{-1}^{\prime}M_{X_{2}(\rho_{n})}\widetilde{U}_{-1}}{E\widetilde{U}_{0}^{2}}\right)^{-1}\frac{n^{-1/2}\widetilde{U}_{-1}^{\prime}M_{X_{2}(\rho_{n})}U}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}=:\nu_{n}\xi_{n},$ (44)
where $\nu_{n}$ and $\xi_{n}$ are defined implicitly by the last equality. Furthermore, we define the martingale difference sequence (MDS)
$\zeta_{i}:=n^{-1/2}\frac{\widetilde{U}_{i-1}U_{i}}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}.$ (45)
We now show that
$n^{-1/2}\frac{\widetilde{U}_{-1}^{\prime}P_{X_{2}(\rho_{n})}U}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}\stackrel{{\scriptstyle p}}{{\to}}0\quad\text{and}\quad\sum_{i=2}^{n}\zeta_{i}\stackrel{{\scriptstyle d}}{{\to}}N(0,1),$ (46)
which establish
$\xi_{n}=\sum_{i=2}^{n}\zeta_{i}-n^{-1/2}\frac{\widetilde{U}_{-1}^{\prime}P_{X_{2}(\rho_{n})}U}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}\stackrel{{\scriptstyle d}}{{\to}}N(0,1).$ (47)
To show the first result in (46), note that
$n^{-1/2}X_{2}(\rho_{n})^{\prime}U=\left(n^{-1/2}\sum_{i=1}^{n}U_{i},n^{-1/2}h_{n}^{1/2}\sum_{i=1}^{n}\rho_{n}^{i-1}U_{i}\right)^{\prime}=(O_{p}(1),O_{p}(1))^{\prime}$ (48)
by the central limit theorem (CLT) for a triangular array of MDS $U_{i}$ and $h_{n}^{1/2}\rho_{n}^{i-1}U_{i}$. Note that
$\frac{h_{n}^{1/2}(1-\rho_{n}^{n})}{n(1-\rho_{n})}$ $=\frac{1}{h_{n}^{1/2}}+o(h_{n}^{-1/2})=o(1)\text{ and}$
$\frac{h_{n}(1-\rho_{n}^{2n})}{n(1-\rho_{n}^{2})}$ $=\frac{1}{2}+o(1).$ (49)
Hence, we have
$(n^{-1}X_{2}(\rho_{n})^{\prime}X_{2}(\rho_{n}))^{-1}=\mathrm{diag}(1,2)+o(1)=O(1).$ (50)
Finally, we have
$n^{-1}(1-\rho_{n})^{1/2}\widetilde{U}_{-1}^{\prime}X_{2}(\rho_{n})=(o_{p}(1),o_{p}(1)),$ (51)
by Lemma 8(a). Then, it follows that
$n^{-1/2}\frac{\widetilde{U}_{-1}^{\prime}P_{X_{2}(\rho_{n})}U}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}$
$=$ $\left(n^{-1/2}X_{2}(\rho_{n})^{\prime}U\right)\left(n^{-1}X_{2}(\rho_{n})^{\prime}X_{2}(\rho_{n})\right)^{-1}\left(n^{-1}\dfrac{\widetilde{U}_{-1}^{\prime}X_{2}(\rho_{n})}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}\right)\stackrel{{\scriptstyle p}}{{\to}}0,$ (52)
where the convergence holds by (48)–(51) and the fact that $E(\widetilde{U}_{0}^{2}U_{1}^{2})=O((1-\rho_{n})^{-1})$ by Lemma 5.
To establish the second result in (46), we adapt the proof of Lemma 1 in [10]. It suffices to verify the analogues of (11) and (12) in [10]; that is, to show the Lindeberg condition
$\sum_{i=2}^{n}E(\zeta_{i}^{2}1(|\zeta_{i}|>\delta)|\mathcal{G}_{i-1})\stackrel{{\scriptstyle p}}{{\to}}0\ \ \text{for any }\delta>0,$ (53)
and
$\sum_{i=2}^{n}E(\zeta_{i}^{2}|\mathcal{G}_{i-1})\stackrel{{\scriptstyle p}}{{\to}}1.$ (54)
The former follows from Lemma 7, while the latter is shown in Lemma 6, using the fact that $\sum_{i=2}^{n}E\zeta_{i}^{2}\to 1.$ Thus, we have proved
$\xi_{n}\stackrel{{\scriptstyle d}}{{\to}}N\left(0,1\right).$ (55)
Next, we show $\nu_{n}\stackrel{{\scriptstyle p}}{{\to}}1$. By Lemma 8(b) and Lemma 5, we have
$\frac{n^{-1}\widetilde{U}_{-1}^{\prime}\widetilde{U}_{-1}}{E\widetilde{U}_{0}^{2}}\stackrel{{\scriptstyle p}}{{\to}}1\quad\text{and}\quad\frac{n^{-1}\widetilde{U}_{-1}^{\prime}P_{X_{2}(\rho_{n})}\widetilde{U}_{-1}}{E\widetilde{U}_{0}^{2}}\stackrel{{\scriptstyle p}}{{\to}}0,$ (56)
which imply $\nu_{n}\stackrel{{\scriptstyle p}}{{\to}}1$. Together with (55), we have
$a_{n}(\widehat{\rho}_{n}(\rho_{n})-\rho_{n})\stackrel{{\scriptstyle d}}{{\to}}N\left(0,1\right).$ (57)
Finally, we show that $d_{n}\widehat{\sigma}_{n}\stackrel{{\scriptstyle p}}{{\to}}1$, which implies that the $t$-statistic $T_{n}\left(\rho_{n}\right)\stackrel{{\scriptstyle d}}{{\to}}N\left(0,1\right)$ for the sub-case $\rho_{n}\to 1$ and $n(1-\rho_{n})\to\infty$. By (56), it is sufficient to show that
$\frac{n^{-1}X_{1}^{\prime}M_{X_{2}(\rho_{n})}\Delta^{2}M_{X_{2}(\rho_{n})}X_{1}}{E(\widetilde{U}_{0}^{2}U_{1}^{2})}\stackrel{{\scriptstyle p}}{{\to}}1.$ (58)
By Lemma 8(e)(g), we have
$(E(\widetilde{U}_{0}^{2}U_{1}^{2}))^{-1}n^{-1}\widetilde{U}_{-1}^{\prime}\Delta^{2}\widetilde{U}_{-1}$ $\stackrel{{\scriptstyle p}}{{\to}}1,$
$(1-\rho_{n})^{1/2}n^{-1}(X_{2}(\rho_{n})^{\prime}\Delta^{2}\widetilde{U}_{-1})$ $=O_{p}(1),\text{ and}$
$n^{-1}(X_{2}(\rho_{n})^{\prime}\Delta^{2}X_{2}(\rho_{n}))$ $=O_{p}(1).$ (59)
These results combined with Lemma 5, (50), and (51) imply (58).
Next, we consider the case where $\rho_{n}\to\rho^{*}<1$. We replicate the steps used in the proof for the case when $n(1-\rho_{n})\to\infty$ and $\rho_{n}\to 1$. Equation (48) holds by the CLT for MDS applied to $U_{i}$ and $h_{n}^{1/2}\rho_{n}^{i-1}U_{i}$, and (49) holds because $h_{n}=n(1-\rho_{n})\to\infty$. Following the same proof strategy as in Lemma 8 in AG12, Lemma 8 also holds when $\rho_{n}\to\rho^{*}<1$. In this case, the bounds involving $1-\rho_{n}$ can be replaced by $1-\rho^{*}$, which is a bounded constant. Therefore, (51) still holds. Note that $E(\widetilde{U}_{0}^{2}U_{1}^{2})=O(1)$ in this case. Hence, we have
$n^{-1/2}\frac{\widetilde{U}_{-1}^{\prime}P_{X_{2}(\rho_{n})}U}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}\stackrel{{\scriptstyle p}}{{\to}}0.$ (60)
Additionally, Lemmas 6 and 7 also hold when $\rho_{n}\to\rho^{*}<1$, hence we have $\sum_{i=1}^{n}\zeta_{i}\stackrel{{\scriptstyle d}}{{\to}}N(0,1)$.
To establish (56) and (58), it suffices to apply the weak law of large numbers (WLLN) for triangular arrays of mean zero, $L^{1+\delta}$ bounded, and $L^{1}$-near-epoch dependent ($L^{1}$-NED) random variables. [6, p.464] shows that the latter conditions imply that the array is a uniformly integrable $L^{1}$ mixingale for which the WLLN holds; see [6, Theorem 2]. For example, we verify that $\widetilde{U}_{i}^{2}-E\widetilde{U}_{0}^{2}$ is $L^{1}$-NED with respect to the $\sigma$-field $\mathcal{G}_{i}$, which implies $n^{-1}\widetilde{U}_{-1}^{\prime}\widetilde{U}_{-1}-E\widetilde{U}_{0}^{2}\stackrel{{\scriptstyle p}}{{\to}}0$ by the WLLN.
By definition of $L^{1}$-NED in [6, p.464], we need to show that there exist constants $\{d_{i}:1\leq i\leq k_{n},n\geq 1\}$ and $\{v_{m}:m\geq 0\}$ such that $v_{m}\downarrow 0$ as $m\to\infty$ and that
$E|\widetilde{U}_{i}^{2}-E(\widetilde{U}_{i}^{2}|\mathcal{G}_{i-m,i+m})|\leq d_{i}v_{m},$ (61)
where $\mathcal{G}_{i-m,i+m}$ is a $\sigma$-filed such that $U_{j}\in\mathcal{G}_{i-m,i+m}$ for $i-m\leq j\leq i+m$. Note that for $m\geq i-1$, we have $E(\widetilde{U}_{i}^{2}|\mathcal{G}_{i-m,i+m})=\widetilde{U}_{i}^{2}$, so that $d_{i}v_{m}=0$. Therefore, we only need to consider the case m<i1m<i-1:
$E(\widetilde{U}_{i}^{2}|\mathcal{G}_{i-m,i+m})$ $=E\left(\left(\sum_{j=i-m}^{i}\rho_{n}^{i-j}U_{j}+\rho_{n}^{m+1}\widetilde{U}_{i-m-1}\right)^{2}|\mathcal{G}_{i-m,i+m}\right)$
$=\left(\sum_{j=i-m}^{i}\rho_{n}^{i-j}U_{j}\right)^{2}+\rho_{n}^{2m+2}E\widetilde{U}_{i-m-1}^{2}.$ (62)
Using (62), we have
$E|\widetilde{U}_{i}^{2}-E(\widetilde{U}_{i}^{2}|\mathcal{G}_{i-m,i+m})|$
[…]
$\leq\$$2\rho_{n}^{2m+2}\sum_{j=1}^{i-m-1}\sum_{k=1}^{i-m-1}\rho_{n}^{2(i-m-1)-j-k}E\left|U_{j}U_{k}\right|+2\rho_{n}^{m+1}\sum_{k=1}^{i-m-1}\sum_{j=i-m}^{i}\rho_{n}^{2i-m-1-j-k}E|U_{j}U_{k}|$
$\leq\$$2M(1-\rho_{n})^{-2}(\rho_{n}^{2m+2}+\rho_{n}^{m+1})\to 0,$ (63)
where the first inequality holds by the triangle inequality and the second inequality holds by $E|U_{j}U_{k}|\leq M$ using the moment condition in part (iv) of the parameter space $\Lambda_{n}$. Therefore, (61) holds by taking $d_{i}=2M(1-\rho_{n})^{-2}$ and $v_{m}=2((\rho^{*})^{2m+2}+(\rho^{*})^{m+1})\downarrow 0$ as $m\to\infty$ because $\rho_{n}\to\rho^{*}<1$.
C.3.  Proofs of Lemmas
Proof of Lemma 1.By a Taylor expansion of $e^{-x}$ around $x=0$, we have
$e^{-x}=1-x+\frac{1}{2}x^{2}-\frac{1}{6}x^{3}+O(x^{4}).$ (64)
Hence,
$e^{-2h}$ $=1-2h+2h^{2}-\frac{4}{3}h^{3}+O(h^{4})\quad\text{and}$
$e^{-hr}$ $=1-hr+\frac{1}{2}h^{2}r^{2}-\frac{1}{6}h^{3}r^{3}+O(h^{4})\text{ as $h\to 0$},$ (65)
where the second equality holds uniformly for $r\in[0,1]$. Substituting these equations into the denominator $h(1-e^{-2h})-2(1-e^{-h})^{2}$ of $\alpha_{h}(r)$ and $\beta_{h}(r)$ gives
$h(1-e^{-2h})-2(1-e^{-h})^{2}=\$$h(2h-2h^{2}+\frac{4}{3}h^{3}+O(h^{4}))-2\left(h-\frac{1}{2}h^{2}+\frac{1}{6}h^{3}+O(h^{4})\right)^{2}$
$=\$$2h^{2}-2h^{3}+\frac{4}{3}h^{4}-2h^{2}\left(1-h+\frac{7}{12}h^{2}\right)+O(h^{5})$
$=\$$\frac{1}{6}h^{4}+O(h^{5})\text{ as $h\to 0$.}$ (66)
Similarly, for the numerators of $\alpha_{h}(r)$ and $\beta_{h}(r)$, we have
$h[1-e^{-2h}-2(1-e^{-h})e^{-hr}+2he^{-hr}-2(1-e^{-h})]$ $=\frac{2-3r}{3}h^{4}+O(h^{5})\quad\text{and}$
$2h[he^{-hr}-(1-e^{-h})]$ $=(2r-1)h^{4}+O(h^{5})\text{ as $h\to 0$}.$ (67)
Therefore, we obtain
$\alpha_{h}(r)$ $=\frac{\frac{2-3r}{3}h^{4}+O(h^{5})}{\frac{1}{6}h^{4}+O(h^{5})}=4-6r+O(h)\quad\text{and}$
$\beta_{h}(r)$ $=\frac{(2r-1)h^{4}+O(h^{5})}{\frac{1}{6}h^{4}+O(h^{5})}=12r-6+O(h)\text{ as $h\to 0$},$ (68)
uniformly over $r\in[0,1]$. Letting $h\downarrow 0$ completes the proof. ∎
Proof of Lemma 2.As $\rho_{n}=1-h_{n}/n$ and $h_{n}\to h\in(0,\infty)$, (69) gives the proofs of parts (a) and (b) of Lemma 2:
$Q$ $=\frac{1-(1-n^{-1}h_{n})^{2n}}{2h_{n}/n-h_{n}^{2}/n^{2}}=\frac{1-e^{-2h_{n}}+o(1)}{2h_{n}/n+o(n^{-1})}=\frac{1-e^{-2h_{n}}}{2h_{n}}n+o(n)\text{ and}$
$S$ $=\frac{1-(1-n^{-1}h_{n})^{n}}{h_{n}/n}=\frac{1-e^{-h_{n}}+o(1)}{h_{n}/n}=\frac{1-e^{-h_{n}}}{h_{n}}n+o(n).$ (69)
Then, we have
$\frac{n^{2}}{nQ-S^{2}}$ $=\frac{n^{2}}{n^{2}(1-e^{-2h_{n}})/2h_{n}-n^{2}(1-e^{-h_{n}})^{2}/h_{n}^{2}+o(n^{2})}$
$=\frac{1}{(1-e^{-2h_{n}})/2h_{n}-(1-e^{-h_{n}})^{2}/h_{n}^{2}}+o(1)$
$\to\frac{2h^{2}}{h(1-e^{-2h})-2(1-e^{-h})^{2}}.$ (70)
Similarly, we compute
$\frac{nS}{nQ-S^{2}}$ $=\frac{n^{2}(1-e^{-h_{n}})/h_{n}+o(n^{2})}{n^{2}(1-e^{-2h_{n}})/2h_{n}-n^{2}(1-e^{-h_{n}})^{2}/h_{n}^{2}+o(n^{2})}$
$=\frac{(1-e^{-h_{n}})/h_{n}}{(1-e^{-2h_{n}})/2h_{n}-(1-e^{-h_{n}})^{2}/h_{n}^{2}}+o(1)$
$\to\frac{2h(1-e^{-h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}$ (71)
and
$\frac{nQ}{nQ-S^{2}}$ $=\frac{n^{2}(1-e^{-2h_{n}})/2h_{n}+o(n^{2})}{n^{2}(1-e^{-2h_{n}})/2h_{n}-n^{2}(1-e^{-h_{n}})^{2}/h_{n}^{2}+o(n^{2})}$
$=\frac{n^{2}(1-e^{-2h_{n}})/2h_{n}}{n^{2}(1-e^{-2h_{n}})/2h_{n}-n^{2}(1-e^{-h_{n}})^{2}/h_{n}^{2}}+o(1)$
$\to\frac{h(1-e^{-2h})}{h(1-e^{-2h})-2(1-e^{-h})^{2}}.$ (72)
For part (f), note that
$n^{-1}\sum_{i=1}^{n}\left(\frac{n(Q-S\rho^{i-1})}{nQ-S^{2}}\right)^{2}$
[…]
$=\$$\left(\frac{nQ}{nQ-S^{2}}\right)^{2}+\left(\frac{nS}{nQ-S^{2}}\right)^{2}n^{-1}Q-2\frac{n^{2}SQ}{(nQ-S^{2})^{2}}n^{-1}S$
[…]
Furthermore, we have
$n^{-1}\sum_{i=1}^{n}\frac{n^{2}(Q-S\rho^{i-1})(n\rho^{i-1}-S)}{(nQ-S^{2})^{2}}$
[…]
$-n^{-1}\sum_{i=1}^{n}\frac{nS}{nQ-S^{2}}\frac{n^{2}}{nQ-S^{2}}\rho_{n}^{2i-2}+n^{-1}\sum_{i=1}^{n}\left(\frac{nS}{nQ-S^{2}}\right)^{2}\rho_{n}^{i-1}$
[…]
$-\frac{nS}{nQ-S^{2}}\frac{n^{2}}{nQ-S^{2}}n^{-1}Q+\left(\frac{nS}{nQ-S^{2}}\right)^{2}n^{-1}S$
[…]
and
$n^{-1}\sum_{i=1}^{n}\left(\frac{n(n\rho^{i-1}-S)}{nQ-S^{2}}\right)^{2}$
[…]
This completes the proof. ∎
Proof of Lemma 3.For part (a), the result holds trivially for $i=1$. When $i=[nr]>1$, letting $h_{n}=n(1-\rho_{n})\to 0$, we have
$\frac{1-\rho_{n}^{i-1}}{n(1-\rho_{n})}=\frac{1-(1-\frac{h_{n}}{n})^{i-1}}{h_{n}}=\frac{1-(1-(i-1)\frac{h_{n}}{n}+{i-1\choose 2}\frac{h_{n}^{2}}{n^{2}})}{h_{n}}+o(1)\to r.$ (76)
For part (b), by Lemma 2(c)(e), we have
$\frac{n(Q-2S+n)}{nQ-S^{2}}$ $=\frac{h_{n}(1-e^{-2h_{n}})-4h_{n}(1-e^{-h_{n}})+2h_{n}^{2}}{h_{n}(1-e^{-2h_{n}})-2(1-e^{-h_{n}})^{2}}+o(1).$ (77)
Using the same Taylor expansion argument as in the proof of Lemma 1, we have
$h_{n}(1-e^{-2h_{n}})-4h_{n}(1-e^{-h_{n}})+2h_{n}^{2}$
[…]
$=\$$\frac{2}{3}h_{n}^{4}+O(h_{n}^{5}),$ (78)
and
$h_{n}(1-e^{-2h_{n}})-2(1-e^{-h_{n}})^{2}=\frac{1}{6}h_{n}^{4}+O(h_{n}^{5}).$ (79)
Therefore, we derive
$\frac{n(Q-2S+n)}{nQ-S^{2}}=\frac{\frac{2}{3}h_{n}^{4}+O(h_{n}^{5})}{\frac{1}{6}h_{n}^{4}+O(h_{n}^{5})}\to 4.$ (80)
For part (c), we have
$\frac{n^{2}(1-\rho_{n})(n-S)}{nQ-S^{2}}$ $=\frac{h_{n}(2h_{n}^{2}-2h_{n}(1-e^{-h_{n}}))}{h_{n}(1-e^{-2h_{n}})-2(1-e^{-h_{n}})^{2}}+o(1)$
$=\frac{2h_{n}^{3}-2h_{n}^{2}(h_{n}-\frac{1}{2}h_{n}^{2}+O(h_{n}^{3}))}{\frac{1}{6}h_{n}^{4}+O(h_{n}^{5})}+o(1)\to 6.$ (81)
For part (d), we have
$\frac{n^{4}(1-\rho_{n})^{2}}{nQ-S^{2}}$ $=\frac{2h_{n}^{4}}{h_{n}(1-e^{-2h_{n}})-2(1-e^{-h_{n}})^{2}}+o(1)$
$=\frac{2h_{n}^{4}}{\frac{1}{6}h_{n}^{4}+O(h_{n}^{5})}+o(1)\to 12.$ (82)
The proof is thus completed. ∎
Proof of Lemma 4.By setting $\phi_{n,i}=1$ in the proof of Lemma 5 in AG12, we have already established this lemma except for parts (e), (g), (i), (l), and certain combinations of part (n). We prove these remaining results below.
For parts (e), (g), (i), and (l), we follow the proof of Lemma 4 in AG12. We prove part (g) here, and the others follow an analogous argument. When $h=0$, we can directly apply Theorem 4.4 of [11]. When $h>0$, we have
$n^{-1/2}\rho_{n}^{[nr]-1}\widetilde{U}_{[nr]}\Rightarrow e^{-hr}\int_{0}^{r}\exp(-(r-s)h)dW(s)=e^{-hr}I_{h}(r),$ (83)
by Theorem 4.4 of [11]. Hence,
$n^{-3/2}\sum_{i=1}^{n}\rho_{n}^{i-1}\widetilde{U}_{n,i-1}\stackrel{{\scriptstyle d}}{{\to}}\int_{0}^{1}e^{-hr}I_{h}(r)dr$ (84)
by the continuous mapping theorem. Parts (e), (i), and (l) can be proved in a similar way and we omit the details.
For part (n), the case with $l_{0}=0$ is established in Lemma 5 of AG12. The cases with $l_{0}=1$ and $2$ follow by the same argument, as the rescaling of $(\rho_{n}^{i-1})^{l_{0}}$ leaves the conclusion unchanged. ∎
Proof of Lemma 8.First, note that
$n^{-1}(1-\rho_{n})^{1/2}\widetilde{U}_{-1}^{\prime}X_{2}(\rho_{n})=\left(n^{-3/2}h_{n}^{1/2}\sum_{i=2}^{n}\widetilde{U}_{i-1},n^{-3/2}h_{n}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right).$ (85)
Also note that $n^{-3/2}h_{n}^{1/2}\sum_{i=1}^{n}\widetilde{U}_{i-1}=o_{p}(1)$ has been proved in Lemma 8 of AG12. For the second term in (85), by Markov’s inequality, it suffices to show that
$n^{-3}h_{n}^{2}E\left(\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)^{2}=o(1).$ (86)
Notice that
$E\left(\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)^{2}$ $=E\left(\sum_{i=2}^{n}\sum_{k=1}^{i-1}\rho_{n}^{i+k-1}U_{i-1-k}\right)^{2}$
$\leq\sum_{i,j=2}^{n}\sum_{k=1}^{i-1}\sum_{l=1}^{j-1}\rho_{n}^{i+j+k+l-2}\left|E(U_{|i+l-(j+k)|+1}U_{1})\right|.$ (87)
By the stationarity of the $U_{i}$’s, the $|E(U_{|i+l-(j+k)|+1}U_{1})|$ term in (87) is uniformly bounded over $i,j,k,l$, and hence, can be ignored.
The contribution of the summands with $i+l=j+k$ is
$\sum_{i=2}^{n}\sum_{k=1}^{i-1}\rho_{n}^{2(i+k)-2}+2\sum_{i>j=2}^{n}\sum_{k=i-j+1}^{i-1}\rho_{n}^{2(j+k)-2}$
[…]
$\leq\$$\left(\sum_{i=2}^{\infty}\rho_{n}^{2i-2}\right)^{2}+(1-\rho_{n})^{-1}\left(\sum_{i=2}^{\infty}\rho_{n}^{i}\right)^{2}=O((1-\rho_{n})^{-3}).$ (88)
This term is negligible because
$(1-\rho_{n})^{-3}n^{-3}h_{n}^{2}=h_{n}^{-1}\to 0.$ (89)
Next, we consider the case $i+l\neq j+k$. By symmetry, we only need to prove it for $i+l>j+k$. We separate this case into two sub-cases (i) l>j+kil>j+k-i, $k\geq i-j+1$ and (ii) l>j+kil>j+k-i, $k\leq i-j$. By the properties of $\alpha$-mixing variables [8, Theorem 3, p. 9] and parts (iii) and (iv) of $\Lambda_{n}$, we have
$E(U_{l}U_{k})\leq 8\|U_{l}\|_{\zeta}\|U_{k}\|_{\zeta}\alpha(|l-k|)^{1-2/\zeta}\leq 8CM|l-k|^{-3(\zeta-2)/(\zeta-3)}$ (90)
for $\zeta>3$. Write $\epsilon=-3/(\zeta-3)$. Hence, the sum for the sub-case (i) can be bounded by
$8CM\sum_{i>j=2}^{n}\sum_{k=1}^{i-1}\sum_{l=j+k-i+1}^{j-1}\rho_{n}^{i+j+k+l-2}(i+l-j-k)^{-3-\epsilon}$
$=\$$8CM\sum_{i>j=2}^{n}\sum_{k=1}^{i-1}\sum_{m=1}^{i-1-k}\rho_{n}^{2(j+k)+m-2}m^{-3-\epsilon}$
$\leq\$$8CM\sum_{i>j=2}^{n}\left(\sum_{k=1}^{i-1}\rho_{n}^{i+j+2k-2}\right)\left(\sum_{m=1}^{i-1-k}\rho_{n}^{m}m^{-3-\epsilon}\right)$
[…]
where the last equality holds by
$\sum_{l=1}^{\infty}\rho_{n}^{2k}=O((1-\rho_{n}^{2})^{-1})=O((1-\rho_{n})^{-1})\text{ and }\sum_{m=1}^{\infty}\rho_{n}^{m}m^{-3-\epsilon}=O(1).$ (92)
Therefore, the summand with $j\neq k$ is also negligible because
$(1-\rho_{n})^{-3}n^{-3}h_{n}^{2}=h_{n}^{-1}\to 0.$ (93)
Similarly, we can prove the sub-case (ii), which is also $O((1-\rho_{n})^{-3})$.
Parts (b) and (c) have been proved in Lemma 8 of AG12.
For part (d), note that
$(n^{-1}X(\rho_{n})^{\prime}X(\rho_{n}))^{-1}n^{-1}X_{2}(\rho_{n})^{\prime}U=(\det)^{-1}(T_{1},T_{2},T_{3})^{\prime},$ (94)
where
$\det:=\$$h_{n}\Bigg[n^{-2}(nQ-S^{2})\left(n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}\right)-n^{-1}\left(n^{-1}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)^{2}$
$+2n^{-1}S\left(n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}\right)\left(n^{-1}\sum_{i=1}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)-n^{-1}Q\left(n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}\right)^{2}\Bigg],$
$T_{1}:=\$$h_{n}\Bigg[\left(n^{-1}Sn^{-1}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}-n^{-1}Qn^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}\right)\left(n^{-1}\sum_{i=1}^{n}U_{i}\right)$
$-\left(n^{-1}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}-n^{-1}Sn^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}\right)\left(n^{-1}\sum_{i=1}^{n}\rho_{n}^{i-1}U_{i}\right)$
$+n^{-2}\left(nQ-S^{2}\right)\left(n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}\right)\Bigg],$
$T_{2}:=\$$h_{n}\Bigg[\left(n^{-1}Qn^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}-\left(n^{-1}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)^{2}\right)\left(n^{-1}\sum_{i=1}^{n}U_{i}\right)$
$-\left(n^{-1}Sn^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}-\left(n^{-1}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)\left(n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}\right)\right)\left(n^{-1}\sum_{i=1}^{n}\rho_{n}^{i-1}U_{i}\right)$
$+n^{-1}(S-h_{n}^{1/2}Q)\left(n^{-1}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)\left(n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}\right)\Bigg],\text{ and}$
$T_{3}:=\$$h_{n}^{1/2}\Bigg[\left(\left(n^{-1}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)\left(n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}\right)-n^{-1}Sn^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}\right)\left(n^{-1}\sum_{i=1}^{n}U_{i}\right)$
$+\left(n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}-\left(n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}\right)^{2}\right)\left(n^{-1}\sum_{i=1}^{n}\rho_{n}^{i-1}U_{i}\right)$
$+\left(n^{-1}Sn^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}-n^{-1}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}\right)\left(n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}\right)\Bigg].$ (95)
Note that
$n^{-2}(nQ-S^{2})$ $=O(n^{-1}(1-\rho_{n})^{-1}),$
$n^{-1}S$ $=O(n^{-1}(1-\rho_{n})^{-1}),$
$n^{-1}Q$ $=O(n^{-1}(1-\rho_{n})^{-1}),$
$n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}$ $=O_{p}((1-\rho_{n})^{-1}),$
$n^{-1}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}$ $=o_{p}(n^{-1/2}(1-\rho_{n})^{-1}),$
$n^{-1}(S-h_{n}^{1/2}Q)$ $=O(n^{-1/2}(1-\rho_{n})^{-1/2}),$
$n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}$ $=O_{p}(n^{-1/2}(1-\rho_{n})^{-1/2}),$
$n^{-1}\sum_{i=1}^{n}\rho_{n}^{i-1}U_{i}$ $=O_{p}(n^{-1}(1-\rho_{n})^{-1/2}),\text{ and }$
$n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}$ $=o_{p}((1-\rho_{n})^{-1/2}),$ (96)
by parts (a) and (b) of Lemma 8, (46), the CLT for MDS for $h_{n}\rho_{n}^{i-1}U_{i}$, and Lemma 5. Using these results, we derive
$\det$ $=O_{p}((1-\rho_{n})^{-2}),$
$T_{1}$ $=O_{p}(n^{-1/2}(1-\rho_{n})^{-1/2}),$
$T_{2}$ $=O_{p}(n^{-1/2}(1-\rho_{n})^{-1}),\text{ and}$
$T_{3}$ $=O_{p}(n^{-1/2}(1-\rho_{n})^{-1}).$ (97)
Therefore, we have
$(\det)^{-1}(T_{1},T_{2},T_{3})^{\prime}=(O_{p}(n^{-1/2}(1-\rho_{n})^{3/2}),O_{p}(n^{-1/2}(1-\rho_{n})),O_{p}(n^{-1/2}(1-\rho_{n})))^{\prime}.$ (98)
For part (e), note that
$\widetilde{U}_{-1}^{\prime}\Delta^{2}\widetilde{U}_{-1}$ $=\sum_{i=2}^{n}\frac{\widetilde{U}_{i-1}^{2}}{(1-p_{ii}^{*})^{2}}\left[U_{i}-(\widetilde{U}_{i-1},1,h_{n}^{1/2}\rho_{n}^{i-1})(X(\rho_{n})^{\prime}X(\rho_{n}))^{-1}X(\rho_{n})^{\prime}U\right]^{2}.$ (99)
By part (c) of the lemma and the fact that $p_{ii}^{*}=O_{p}(n^{-1/2})$, it remains to be shown that
$O((1-\rho_{n}))n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}U_{i}(\widetilde{U}_{i-1},1,h_{n}^{1/2}\rho_{n}^{i-1})(X(\rho_{n})^{\prime}X(\rho_{n}))^{-1}X(\rho_{n})^{\prime}U$ $\stackrel{{\scriptstyle p}}{{\to}}0\text{ and}$
$O((1-\rho_{n}))n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}\left[(\widetilde{U}_{i-1},1,h_{n}^{1/2}\rho_{n}^{i-1})(X(\rho_{n})^{\prime}X(\rho_{n}))^{-1}X(\rho_{n})^{\prime}U\right]^{2}$ $\stackrel{{\scriptstyle p}}{{\to}}0.$ (100)
By part (d), we only need to show that
$(1-\rho_{n})^{2}n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}U_{i}$ $=o_{p}(1),$
$(1-\rho_{n})^{2}n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{4}$ $=o_{p}(1),$
$(1-\rho_{n})^{2}n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}U_{i}$ $=o_{p}(1),$
$(1-\rho_{n})^{3}n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}$ $=o_{p}(1),$
$(1-\rho_{n})^{2}n^{-3/2}\sum_{i=2}^{n}h_{n}^{1/2}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}U_{i}$ $=o_{p}(1),$
$(1-\rho_{n})^{3}n^{-2}\sum_{i=2}^{n}h_{n}^{1/2}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}$ $=o_{p}(1),$
$(1-\rho_{n})^{3}n^{-2}\sum_{i=2}^{n}h_{n}\rho_{n}^{2i-2}\widetilde{U}_{i-1}^{2}$ $=o_{p}(1),\text{ and}$
$(1-\rho_{n})^{5/2}n^{-2}\sum_{i=2}^{n}h_{n}^{1/2}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}$ $=o_{p}(1).$ (101)
We only need to prove the last four equations in (101) because all the other results have been established in Lemma 8 of AG12. Using the facts that
$n^{-1}\sum_{i=2}^{n}h_{n}^{1/2}\rho_{n}^{i-1}\to 0\text{ and }n^{-1}\sum_{i=2}^{n}h_{n}\rho_{n}^{2i-2}\to 1/2,$ (102)
the sixth and seventh lines are proved by part (b) of the lemma. The fifth and last lines of (101) can be proved using Markov’s inequality, as in the proof of part (a), together with the methods employed in the proof of Lemma 7 in AG12.
For part (f), we have
$\widetilde{U}_{-1}^{\prime}\Delta^{2}X_{2}(\rho_{n})$
[…]
Then, the desired results follow from
$(1-\rho_{n})^{1/2}n^{-1}\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}^{2}(1,h_{n}\rho_{n}^{i-1})$ $=(O_{p}(1),O_{p}(1))\text{ and}$
$(1-\rho_{n})^{5/2}n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}\bar{U}_{i-1}^{2}(1,h_{n}\rho_{n}^{i-1})$ $=(O_{p}(1),O_{p}(1)),$ (104)
where $\bar{U}_{i-1}=(\widetilde{U}_{i-1},1,h_{n}^{1/2}\rho_{n}^{i-1})(X(\rho_{n})^{\prime}X(\rho_{n}))^{-1}X(\rho_{n})^{\prime}U$.
To prove the first line of (104), by Markov’s inequality we only need to show that
$(1-\rho_{n})n^{-2}E\left[(\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}^{2})^{2}\right]$ $=O(1)\text{ and}$
$(1-\rho_{n})h_{n}^{2}n^{-2}E\left[(\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}U_{i}^{2})^{2}\right]$ $=O(1).$ (105)
Note that
$E\left(\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}^{2}\right)^{2}$
[…]
$+64C^{2}M\sum_{i,j=2}^{n}\sum_{k=1}^{i-1}\sum_{l=1}^{j-1}\rho_{n}^{k+l}(k+1)^{-3-\epsilon}(l+1)^{-3-\epsilon}$
[…]
$+64C^{2}Mn^{2}\left(\sum_{k=1}^{\infty}(k+1)^{-3-\epsilon}\right)^{2}$
[…]
where the second inequality uses the properties of $\alpha$-mixing analogously to (90) and the moment conditions in parts (iii) and (iv) of $\Lambda_{n}$, and the last equality holds by
$\sum_{i=1}^{\infty}i^{-3-\epsilon}$ $=O(1)\text{ and }\sum_{k=1}^{\infty}\rho_{n}^{k}=O((1-\rho_{n})^{-1}).$ (107)
Therefore, we have
$(1-\rho_{n})n^{-2}E\left[(\sum_{i=2}^{n}\widetilde{U}_{i-1}U_{i}^{2})^{2}\right]=O(h_{n}^{-1}+n^{-1}h_{n})=o(1),$ (108)
which establishes the first line of (105).
For the second line of (105), using the same expansion and results as before, we have
$E(\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}U_{i}^{2})^{2}\leq\$$M\sum_{i=j\geq 2}^{n}\rho_{n}^{i+j-2}\sum_{k=1}^{i-1}\sum_{l=1}^{j-1}\rho_{n}^{k+l}$
$+8CM\sum_{i,j\geq 2,i\neq j}\rho_{n}^{i+j-2}\sum_{k=1}^{i-1}\sum_{l=1}^{j-1}\rho_{n}^{k+l}|i-j|^{-3-\epsilon}$
$+64C^{2}M\sum_{i,j=2}^{n}\rho_{n}^{i+j-2}\sum_{k=1}^{i-1}\sum_{l=1}^{j-1}\rho_{n}^{k+l}(k+1)^{-3-\epsilon}(l+1)^{-3-\epsilon}$
[…]
$+8CM\left(\sum_{i,j\geq 2,i\neq j}^{\infty}\rho_{n}^{i+j-2}|i-j|^{-3-\epsilon}\right)\left(\sum_{k=1}^{\infty}\rho_{n}^{k}\right)^{2}$
$+64C^{2}M\left(\sum_{k=1}^{\infty}\rho_{n}^{k}\right)^{2}\left(\sum_{k=1}^{\infty}(k+1)^{-3-\epsilon}\right)^{2}$
[…]
where the last equality holds by $\sum_{k=1}^{\infty}\rho_{n}^{2k-2}=O((1-\rho_{n})^{-1})$ and
$\sum_{i,j\geq 2,i\neq j}^{\infty}\rho_{n}^{i+j-2}|i-j|^{-3-\epsilon}$ $=2\sum_{i>j\geq 2}^{\infty}\rho_{n}^{i+j-2}(i-j)^{-3-\epsilon}=2\sum_{j=2}^{\infty}\sum_{k=1}^{\infty}\rho_{n}^{k+2j-2}k^{-3-\epsilon}$
$=2\left(\sum_{j=2}^{\infty}\rho_{n}^{2j-2}\right)\left(\sum_{k=1}^{\infty}\rho_{n}^{k}k^{-3-\epsilon}\right)=O((1-\rho_{n})^{-1}).$ (110)
Therefore, we have
$(1-\rho_{n})h_{n}^{2}n^{-2}E\left[(\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}U_{i}^{2})^{2}\right]=O((1-\rho_{n})^{-2}h_{n}^{2}n^{-2})=O(1).$ (111)
To prove the second line of (104), note that
$\bar{U}_{i-1}^{2}=\$$O_{p}(n^{-1}(1-\rho_{n})^{2})[1+h_{n}\rho_{n}^{2i-2}+(1-\rho_{n})\widetilde{U}_{i-1}^{2}$
$+2h_{n}^{1/2}\rho_{n}^{i-1}+2(1-\rho_{n})^{1/2}\widetilde{U}_{i-1}+2h_{n}^{1/2}(1-\rho_{n})\rho_{n}^{i-1}\widetilde{U}_{i-1}]$ (112)
by part (d). Following the same argument as in part (a), we can show that
$(1-\rho_{n})^{5/2}n^{-2}h_{n}^{k_{0}}\sum_{i=2}^{n}\rho_{n}^{k_{1}(i-1)}\widetilde{U}_{i-1}$ $=O_{p}(1)$ (113)
for $(k_{0},k_{1})=(0,0),(1/2,1),(1,2),(1,1),(3/2,2)$, and $(2,3)$. For example, the case $(2,3)$ can be established by proving
$(1-\rho_{n})^{5}n^{-4}h_{n}^{4}E\left(\sum_{i=2}^{n}\rho_{n}^{3(i-1)}\widetilde{U}_{i-1}\right)^{2}=(1-\rho_{n})^{5}n^{-4}h_{n}^{4}\times O((1-\rho_{n})^{-3})=o(1),$ (114)
following the same expansion as in (87) and observing that $\sum_{j=1}^{n}\rho_{n}^{6j}=O((1-\rho_{n})^{-1})$. Then we only need to show that
$(1-\rho_{n})^{l_{0}}n^{-2}h_{n}^{l_{1}}\sum_{i=2}^{n}\rho_{n}^{l_{2}(i-1)}\widetilde{U}_{i-1}^{l_{3}}$ $=O_{p}(1),$ (115)
for $(l_{0},l_{1},l_{2},l_{3})=(7/2,0,0,3)$, $(3,0,0,2)$, $(7/2,1/2,1,2)$, $(7/2,1,1,3)$, $(3,1,1,2)$, and (7/2, 3/2,2,2).
The cases $(3,0,0,2)$ and $(7/2,1/2,1,2)$ are implied by the fourth and sixth lines of (101), respectively.
For the case $(7/2,0,0,3)$, we have
$\left|(1-\rho_{n})^{7/2}n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}\right|$ $\leq(1-\rho_{n})^{2}\left((1-\rho_{n})^{2}n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{4}\right)^{1/2}\left((1-\rho_{n})n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}\right)^{1/2}$
$=(1-\rho_{n})^{2}\times o_{p}(1)\times O_{p}(n^{-1/2})=o_{p}(1),$ (116)
by the Cauchy-Schwarz inequality, the second line of (101), part (b), and Lemma 5.
For the case $(7/2,3/2,2,2)$, note that $E\widetilde{U}_{i-1}^{2}=O(E\widetilde{U}_{0}^{2})=O((1-\rho_{n})^{-1})$ by Lemma 5. Hence, we have
$(1-\rho_{n})^{7/2}h_{n}^{3/2}n^{-2}E\left[\sum_{i=2}^{n}\rho_{n}^{2i-2}\widetilde{U}_{i-1}^{2}\right]$ $=(1-\rho_{n})^{7/2}h_{n}^{3/2}n^{-2}\sum_{i=2}^{n}\rho_{n}^{2i-2}O((1-\rho_{n})^{-1}$
$=O((1-\rho_{n})^{3}n^{-1/2})=o(1).$ (117)
The result follows by Markov’s inequality.
For the case $(7/2,1,1,3)$, we use the Cauchy-Schwarz inequality:
$(1-\rho_{n})^{7/2}n^{-2}h_{n}\left|\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}\right|$
$\leq\$$\left((1-\rho_{n})^{7/2}h_{n}^{2}n^{-2}\sum_{i=2}^{n}\rho_{n}^{2i-2}\widetilde{U}_{i-1}^{2}\right)^{1/2}\left((1-\rho_{n})^{7/2}n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{4}\right)^{1/2}$
[…]
by the second line of (101) and (117). For the case $(3,1,1,2)$, by the Cauchy-Schwarz inequality, we have
$(1-\rho_{n})^{3}h_{n}\left|n^{-2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}\right|$
$\leq\$$(1-\rho_{n})h_{n}^{1/2}\left(n^{-2}(1-\rho_{n})^{3}\sum_{i=2}^{n}h_{n}\rho_{n}^{2i-2}\widetilde{U}_{i-1}^{2}\right)^{1/2}\left(n^{-2}(1-\rho_{n})\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}\right)^{1/2}$
$=\$$(1-\rho_{n})h_{n}^{1/2}\times o_{p}(1)\times O_{p}(n^{-1/2})=o_{p}(1),$ (119)
by the seventh line of (101), part (b), and Lemma 5.
Finally, for part (g), note that
$X_{2}(\rho_{n})^{\prime}\Delta^{2}X_{2}(\rho_{n})=$ $\sum_{i=1}^{n}\frac{(1+2h_{n}^{1/2}\rho_{n}^{i-1}+h_{n}\rho_{n}^{2i-2})}{(1-p_{ii}^{*})^{2}}\left(U_{i}-\bar{U}_{i-1}\right)^{2}.$ (120)
As $p_{ii}^{*}=O_{p}(n^{-1/2})$, we only need to establish the following:
$n^{-1}\sum_{i=1}^{n}\left(1+2h_{n}^{1/2}\rho_{n}^{i-1}+h_{n}\rho_{n}^{2i-2}\right)U_{i}^{2}$ $=O_{p}(1)\text{ and}$
$n^{-1}\sum_{i=1}^{n}\left(1+2h_{n}^{1/2}\rho_{n}^{i-1}+h_{n}\rho_{n}^{2i-2}\right)\bar{U}_{i-1}^{2}$ $=O_{p}(1).$ (121)
Because $E[U_{i}^{2}]=O(1)$, the first line holds by
$n^{-1}\sum_{i=1}^{n}\left(1+2h_{n}^{1/2}\rho_{n}^{i-1}+h_{n}\rho_{n}^{2i-2}\right)\leq 1+2h_{n}^{1/2}n^{-1}(1-\rho_{n})^{-1}+h_{n}n^{-1}(1-\rho_{n}^{2})^{-1}=O(1).$ (122)
By (112), it remains to show that
$n^{-2}(1-\rho_{n})^{2}\sum_{i=1}^{n}\left(1+2h_{n}^{1/2}\rho_{n}^{i-1}+h_{n}\rho_{n}^{2i-2}\right)^{2}$ $=O(1),$
$n^{-2}(1-\rho_{n})^{3}\sum_{i=2}^{n}\left(1+2h_{n}^{1/2}\rho_{n}^{i-1}+h_{n}\rho_{n}^{2i-2}\right)\widetilde{U}_{i-1}^{2}$ $=O_{p}(1),\text{ and}$
$n^{-2}(1-\rho_{n})^{5/2}\sum_{i=2}^{n}\left(1+2h_{n}^{1/2}\rho_{n}^{i-1}+h_{n}\rho_{n}^{2i-2}\right)\left(1+h_{n}^{1/2}(1-\rho_{n})^{1/2}\rho_{n}^{i-1}\right)\widetilde{U}_{i-1}$ $=O_{p}(1).$ (123)
The first equality holds by (122) and the fact that
$n^{-2}(1-\rho_{n})^{2}h_{n}^{3/2}\sum_{i=1}^{n}\rho_{n}^{3i-3}=o(1)\text{ and }n^{-2}(1-\rho_{n})^{2}h_{n}^{2}\sum_{i=1}^{n}\rho_{n}^{4i-4}=o(1).$ (124)
By Markov’s inequality, the second equality holds by (122) and the fact $E\widetilde{U}_{i-1}^{2}=O((1-\rho_{n})^{-1})$. Finally, the last equality holds by part (a) and (113) for cases $(1,2)$ and $(2,3)$. ∎
Appendix D Additional Simulation Results
We provide additional simulation results in this section. Table SM.1 reports the CPs of [13, Mik07 hereafter] CIs. [^4] The Mik07 CI differs from the [4] CI in that there is no initial-condition component in the asymptotic distribution of the test statistic, and a homoskedastic variance estimator is used. The CPs of Mik07 CIs are close to the nominal level when the $U_{i}$ sequences are i.i.d. and $Y_{0}^{*}$ is either fixed or stationary. When $Y_{0}^{*}$ is drawn from scaled $n$ or explosive distributions, or when the $U_{i}$ sequences are non-i.i.d., the CPs of Mik07 CIs fall below 95%.
Table SM.1: Coverage probabilities ($\times 100$) of nominal 95% Mik07 CIs
Initial Conditions: Fixed Stationary
$\bm{\rho}$: .00 .50 .70 .90 .99 .00 .50 .70 .90 .99
i.i.d. 94.9 94.7 94.9 94.8 94.7 94.8 94.7 94.9 94.7 94.0
GARCH1 93.8 93.8 94.3 94.4 94.5 93.8 93.8 94.2 94.4 93.6
GARCH2 89.6 90.6 90.7 92.3 93.4 89.6 90.7 90.6 92.4 92.8
GARCH3 85.3 86.0 87.3 89.4 92.3 85.3 86.0 87.1 89.6 91.8
ARCH4 80.3 81.6 82.5 86.3 91.1 80.4 81.6 82.7 86.5 90.8
Initial Conditions: Scaled $\bm{n}$ Explosive
i.i.d. 94.8 94.6 94.7 93.5 80.4 94.6 94.4 94.2 92.3 73.5
GARCH1 94.3 93.8 93.7 92.7 79.8 94.3 93.8 93.6 91.5 73.0
GARCH2 91.1 91.9 91.7 92.2 80.3 92.4 92.8 92.5 91.4 73.0
GARCH3 87.5 88.5 89.2 90.8 80.2 90.2 90.7 91.2 90.9 73.4
ARCH4 84.0 84.8 86.5 89.2 80.4 88.5 89.0 89.7 90.3 73.6
Table SM.2 reports the ratios of the ALs of ICR to Mik07 CIs when the $U_{i}$ sequences are i.i.d. and $Y_{0}^{*}$ is fixed or stationary. The ratios are between 1 and 1.12, suggesting that the ALs of ICR CIs are similar to those of Mik07 CIs when the true data generating process is such that the Mik07 inference method is valid. Tables SM.3 and SM.4 report the ALs of [4, AG14 hereafter] and Mik07 CIs, respectively. The results are consistent with the conclusions drawn from Tables 4 and SM.2.
Table SM.2: Ratios of average lengths of nominal 95% ICR to Mik07 CIs
Initial Conditions: Fixed Stationary
$\bm{\rho}$: .00 .50 .70 .90 .99 .00 .50 .70 .90 .99
i.i.d. 1.00 1.01 1.02 1.03 1.09 1.01 1.01 1.03 1.05 1.12
Table SM.3: Average lengths of nominal 95% AG14 CIs
Initial Conditions: Fixed Stationary
$\bm{\rho}$: .00 .50 .70 .90 .99 .00 .50 .70 .90 .99
i.i.d. .32 .28 .24 .16 .07 .32 .28 .24 .16 .07
GARCH1 .32 .28 .24 .16 .07 .33 .29 .25 .17 .07
GARCH2 .38 .34 .28 .18 .08 .38 .33 .28 .18 .07
GARCH3 .43 .38 .32 .19 .08 .43 .38 .32 .19 .07
ARCH4 .49 .43 .36 .21 .09 .48 .43 .36 .20 .08
Initial Conditions: Scaled $\bm{n}$ Explosive
i.i.d. .22 .19 .16 .09 .02 .10 .09 .08 .04 .01
GARCH1 .05 .06 .07 .05 .02 .00 .01 .02 .02 .01
GARCH2 .27 .23 .19 .10 .02 .13 .12 .09 .04 .01
GARCH3 .31 .27 .21 .11 .02 .15 .14 .11 .05 .01
ARCH4 .35 .30 .24 .12 .03 .17 .15 .12 .05 .01
Table SM.4: Average lengths of nominal 95% Mik07 CIs
Initial Conditions: Fixed Stationary
$\bm{\rho}$: .00 .50 .70 .90 .99 .00 .50 .70 .90 .99
i.i.d. .32 .28 .24 .17 .07 .32 .28 .24 .16 .07
GARCH1 .31 .27 .23 .16 .07 .32 .28 .24 .16 .07
GARCH2 .32 .28 .24 .17 .07 .32 .28 .24 .16 .07
GARCH3 .32 .28 .24 .17 .08 .32 .28 .24 .16 .07
ARCH4 .32 .28 .24 .17 .08 .32 .28 .24 .16 .07
Initial Conditions: Scaled $\bm{n}$ Explosive
i.i.d. .24 .20 .16 .09 .02 .14 .11 .08 .04 .01
GARCH1 .10 .09 .08 .05 .02 .02 .02 .02 .02 .01
GARCH2 .25 .21 .17 .09 .02 .15 .12 .09 .04 .01
GARCH3 .25 .21 .17 .10 .02 .15 .12 .09 .05 .01
ARCH4 .25 .21 .17 .10 .02 .15 .12 .10 .05 .01

### v3

@@
 We consider an AR model with an AR parameter that may take values close or equal to one.
-Existing CIs in the literature assume that the initial condition is stationary, zero, or fixed, e.g., see [14], [7], [1], [12], [9], [13], and [4, AG14 hereafter].
-If the initial condition assumption is not satisfied, then these CIs do not have correct asymptotic coverage probabilities and their finite sample coverage probabilities can be poor.
+Existing CIs in the literature for the AR parameter assume that the initial condition is stationary, zero, or fixed, e.g., see [15], [7], [1], [12], [9], [14], and [4, AG14 hereafter].
+If the initial-condition assumption is not satisfied, then these CIs do not have correct asymptotic coverage probabilities and their finite sample coverage probabilities can be poor.
 For example, simulations reported below show that the nominal 95% AG14 CI has finite sample coverage probabilities ranging from 24.1% to 93.5% across 50 cases with highly variable initial conditions and different types of conditional heteroskedasticity of the error.
@@
 These CIs are easy to compute and do not depend on any tuning parameters.
-We show that the ICR CI has correct asymptotic size in a uniform sense under conditions that allow for arbitrary initial condition and conditional heteroskedasticity.
+We show that the ICR CI has correct asymptotic size in a uniform sense under conditions that allow for an arbitrary initial condition and conditional heteroskedasticity.
 Simulations show that the finite sample CPs of the nominal 95% ICR CI are quite good.
@@
 However, simulations show that the effect is fairly small.
-Across 50 scenarios with stationary or fixed initial conditions and i.i.d., GARCH, or ARCH errors, the ratio of the expected length of the ICR CI compared to that of the AG14 CI was found to vary between 1.00 and 1.11.
-On average, the ICR CIs are 3.5% longer than the AG14 CIs in these scenarios.
+Across 50 scenarios with stationary or fixed initial conditions and i.i.d., GARCH, or ARCH errors, the ratio of the expected length of the ICR CI compared to that of the AG14 CI is found to vary between 1.00 and 1.11.
+On average, the ICR CI is 3.5% longer than the AG14 CI in these scenarios.
 Based on the ICR CI, we also introduce an asymptotically median-unbiased interval estimator (MUE) of the AR parameter.
+The AR model considered here has been applied in the literature to exchange rate, commodity and stock prices, and other economic time series, e.g., see [13].
 This paper is organized as follows.
@@
 We consider the AR(1) model with conditional-heteroskedasticity studied in AG14:
-$Y_{i}$ $=\mu+Y_{i}^{*}\quad\text{and}$
+$Y_{i}$ $=\mu+Y_{i}^{*}\quad\text{for $i=0,1,\dots,n$ and}$
 $Y_{i}^{*}$ $=\rho Y_{i-1}^{*}+U_{i}\quad\text{for }i=1,2,\dots,n,$ (1)
-where $\rho\in[-1+\varepsilon,1]$ for some $0<\varepsilon<2$ and $\{U_{i}:i=1,2,\dots\}$ are stationary and ergodic under the distribution $F$, with conditional mean 0 given a $\sigma$-field $\mathcal{G}_{i-1}$ for which $U_{j}\in\mathcal{G}_{i}$ for all $j\leq i$, conditional variance $\sigma_{i}^{2}=E_{F}(U_{i}^{2}|\mathcal{G}_{i-1})$, and unconditional variance $\sigma_{U}^{2}\in(0,\infty)$.
+where the observed data are $\{Y_{0},\dots,Y_{n}\}$, $\rho\in[-1+\varepsilon,1]$ for some $0<\varepsilon<2$, and $\{U_{i}:i=1,2,\dots\}$ are stationary and ergodic under the distribution $F$, with conditional mean 0 given a $\sigma$-field $\mathcal{G}_{i-1}$ for which $U_{j}\in\mathcal{G}_{i}$ for all $j\leq i$, conditional variance $\sigma_{i}^{2}=E_{F}(U_{i}^{2}|\mathcal{G}_{i-1})$, and unconditional variance $\sigma_{U}^{2}\in(0,\infty)$.
 A detailed description of the parameter space $\Lambda$ is provided in Section 5.1.
 By iterative substitution in (1), we have
-$Y_{i}$ $=\mu+\rho^{i-1}\cdot\rho\cdot Y_{0}^{*}+\sum_{j=1}^{n}\rho^{i-j}U_{j}\quad\text{for }i=1,\dots,n.$ (2)
+$Y_{i}$ $=\mu+\rho^{i-1}\cdot\rho\cdot Y_{0}^{*}+\sum_{j=1}^{i}\rho^{i-j}U_{j}\quad\text{for }i=1,\dots,n.$ (2)
 Let $Y$, $U$, and $X_{1}$ be $n$-vectors whose $i$-th elements are $Y_{i}$, $U_{i}$, and $Y_{i-1}$, respectively.
@@
 We compare the coverage probabilities (CPs) and average lengths (ALs) of the ICR and AG14 CIs and report the absolute median biases (ABMs) of the ICR MUE using Monte Carlo simulations.
-In Section D of the Supplemental Material, we compare the performance of the ICR and [13, Mik07 hereafter] CIs.
+In Section D of the Supplemental Material, we compare the performance of the ICR and [14, Mik07 hereafter] CIs.
 We focus on the nominal 95% equal-tailed two-sided CIs.
@@
 Table 3 shows that the ICR CI achieves CPs close to 95% in all of the scenarios considered and for any initial conditions.
+Specifically, all ICR CI CPs lie in the interval [93.5, 95.0].
 Table 4 reports the ratios of the ALs of the nominal 95% ICR CI to the AG14 CI.
-We include only cases with fixed or stationary initial conditions because AG14 CIs are not robust to more variable initial conditions, whereas ICR CIs are.
+We include only cases with fixed or stationary initial conditions because the AG14 CI is not robust to more variable initial conditions, whereas the ICR CI is.
 The ratios are between 1.00 and 1.11.
-On average, the ICR CIs are 3.5% longer than the AG14 CIs over the fifty cases considered.
+On average, the ICR CI is 3.5% longer than the AG14 CI over the fifty cases considered.
 Thus, the price the ICR CI pays in terms of AL to gain robustness against the distribution of the initial condition is fairly small.
-Table 2: Coverage probabilities ($\times 100$) of nominal 95% AG14 CIs
+Table 2: Coverage probabilities ($\times 100$) of the nominal 95% AG14 CI
 Initial Conditions: Fixed Stationary
@@
 ARCH4 91.1 92.3 92.8 93.5 84.4 76.7 86.3 89.4 91.8 77.9
-Table 3: Coverage probabilities ($\times 100$) of nominal 95% ICR CIs
+Table 3: Coverage probabilities ($\times 100$) of the nominal 95% ICR CI
 Initial Conditions: Arbitrary
@@
 ARCH4 93.5 93.8 93.9 94.5 94.3
-Table 4: Ratios of average lengths of nominal 95% ICR to AG14 CIs
+Table 4: Ratios of the average lengths of the nominal 95% ICR to AG14 CIs
 Initial Conditions: Fixed Stationary
@@
 ARCH4 1.00 1.02 1.03 1.01 1.06 1.01 1.02 1.04 1.03 1.10
-Table 5 presents the ALs of ICR CIs. These are substantially decreasing in $\rho$, as expected, longer for ARCH than i.i.d., and increasing from GARCH1 to GARCH2 to GARCH3.
-Table 5: Average lengths of nominal 95% ICR CIs
+Table 5: Average lengths of the nominal 95% ICR CI
 Initial Conditions: Fixed Stationary
@@
 ARCH4 .46 .40 .32 .17 .04 .39 .32 .25 .12 .02
-Table 6 reports the AMBs of the ICR MUE. Across all cases, the AMBs range from 0.000 to 0.022, indicating that their magnitudes are generally small.
-Table 6: Absolute Median Biases of ICR MUEs
+Table 5 presents the ALs of the ICR CI.
+These are substantially decreasing in $\rho$, as expected, longer for ARCH conditional heteroskedasticity than i.i.d., and increasing from GARCH1 to GARCH2 to GARCH3 conditional heteroskedasticity.
+Table 6: Absolute Median Biases of the ICR MUE
 Initial Conditions: Fixed Stationary
@@
 ARCH4 .022 .016 .011 .005 .010 .022 .016 .011 .010 .010
+Table 6 reports the AMBs of the ICR MUE. Across all cases, the AMBs range from 0.000 to 0.022, indicating that their magnitudes are generally small.
 In Section D of the Supplemental Material, we compare the performance of the ICR and Mik07 CIs via simulations.
@@
 yapengzheng@link.cuhk.edu.hk.
-February 7, 2026
+March 11, 2026
 In this Supplemental Material, Section A describes how the critical values are computed, Section B proves Theorem 1, Section C proves Theorem 2, and Section D provides some additional simulation results.
@@
 We initialize the process at $W(0)=0$ and construct each path iteratively as
-$W(r_{i})=W(r_{i-1})+Z_{i-1}/N^{1/2},\ \ r_{i}=i/N,\ i=1,...,N,$ (14)
-where $Z_{i-1}$ are independent standard normal random variables. Next, we compute $I_{h}(r_{j})$, $j=1,\dots,N$, using (10) via a numerical approximation:
+$W(r_{i})=W(r_{i-1})+Z_{i-1}/N^{1/2}\ \text{ for }r_{i}=i/N\text{ and }i=1,...,N,$ (14)
+where $Z_{i-1}$ are independent standard normal random variables. Next, we compute $I_{h}(r_{j})$ for $j=1,\dots,N$, using (10) via a numerical approximation:
 $I_{h}(r_{j})=\begin{cases}\sum_{i=1}^{j}\exp(-(r_{j}-r_{i})h)\cdot(W(r_{i})-W(r_{i-1}))&\text{for }h>0,\\ W(r_{j})&\text{for }h=0.\end{cases}$ (15)
-Then, we use (10) to compute $I_{f,h}(r_{j})$, $j=1,\dots,N$.
+Then, we use (10) to compute $I_{f,h}(r_{j})$ for $j=1,\dots,N$.
 Given $I_{f,h}(r_{j})$ and $W(r_{j})$, we evaluate $J_{h}$ in (9) via numerical approximation.
@@
 Lemma 2.
-When $n(1-\rho_{n})\to h\in(0,\infty)$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
+When $n(1-\rho_{n})\to h\in(0,\infty)$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly:
   1. (a)
@@
 Lemma 3.
-When $n(1-\rho_{n})\to h=0$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
+When $n(1-\rho_{n})\to h=0$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly:
   1. (a)
@@
 Lemma 4.
-When $n(1-\rho_{n})\to h\in(0,\infty)$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
+When $n(1-\rho_{n})\to h\in(0,\infty)$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly:
   1. (a)
@@
 Lemma 5.
-When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly
+When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly:
 $E(\widetilde{U}_{0}^{2}U_{1}^{2})-(1-\rho_{n}^{2})^{-1}(EU_{1}^{2})^{2}$ $=O(1),$
@@
 Lemma 8.
-When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly, we have
+When $n(1-\rho_{n})\to h=\infty$, for a sequence $\{\lambda_{n}=(\rho_{n},F_{n})\in\Lambda_{n}\}_{n\geq 1}$, the following results hold jointly:
   1. (a)
@@
 $-\frac{2h^{2}}{h(1-e^{-2h})-2(1-e^{-h})^{2}}\left(\int_{0}^{1}e^{-hr}I_{h}(r)dr\right)^{2}$
-$=\$$\int_{0}^{1}I_{f,h}^{2}(r)dr.$ (21)
+$=\$$\int_{0}^{1}I_{f,h}^{2}(r)dr,$ (21)
 where the convergence holds by Lemma 2 and parts (f) and (g) of Lemma 4 and the last equality holds by (10).
@@
 Since $P(\int_{0}^{1}I_{f,h}^{2}(r)dr=0)=0$, we have $P(\det(B_{n})=0)\to 0$ as $n\to\infty$, which implies that $B_{n}$ is invertible with probability approaching one.
-Combining it with the fact that $B_{n}=O_{p}(1)$, we have $B_{n}^{-1}=O_{p}(1)$.
+Combining this with the fact that $B_{n}=O_{p}(1)$, we have $B_{n}^{-1}=O_{p}(1)$.
 Next, we define
 $D_{n}$ $=\begin{pmatrix}n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}U_{i}\\ n^{-3/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}\rho_{n}^{i-1}U_{i}\\ n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}U_{i}\end{pmatrix}\text{ and}$
-$K_{n}$ $=\begin{pmatrix}n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}&n^{-2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}&n^{-5/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}\\ n^{-2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}&n^{-2}\sum_{i=2}^{n}\rho_{n}^{2i-2}\widetilde{U}_{i-1}^{2}&n^{-5/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}\\ n^{-5/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}&n^{-5/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}&n^{-3}\sum_{i=2}^{n}\widetilde{U}_{j-1}^{4}\end{pmatrix}$.
+$K_{n}$ $=\begin{pmatrix}n^{-2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{2}&n^{-2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}&n^{-5/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}\\ n^{-2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{2}&n^{-2}\sum_{i=2}^{n}\rho_{n}^{2i-2}\widetilde{U}_{i-1}^{2}&n^{-5/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}\\ n^{-5/2}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{3}&n^{-5/2}\sum_{i=2}^{n}\rho_{n}^{i-1}\widetilde{U}_{i-1}^{3}&n^{-3}\sum_{i=2}^{n}\widetilde{U}_{i-1}^{4}\end{pmatrix}$.
 (28)
@@
 […]
-$+n^{-2}\frac{1}{(nQ-S^{2})^{2}}\sum_{i=1}^{n}\left(\sum_{k=2}^{n}\left(Q-S(\rho_{n}^{k-1}+\rho_{n}^{i-1})+n\rho^{i+k-2}\right)\widetilde{U}_{k-1}\right)^{2}\widehat{U}_{i}^{2}+o_{p}(1)$
+$+n^{-2}\frac{1}{(nQ-S^{2})^{2}}\sum_{i=1}^{n}\left(\sum_{k=2}^{n}\left(Q-S(\rho_{n}^{k-1}+\rho_{n}^{i-1})+n\rho_{n}^{i+k-2}\right)\widetilde{U}_{k-1}\right)^{2}\widehat{U}_{i}^{2}+o_{p}(1)$
 […]
@@
 $n^{-1/2}\frac{\widetilde{U}_{-1}^{\prime}P_{X_{2}(\rho_{n})}U}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}\stackrel{{\scriptstyle p}}{{\to}}0\quad\text{and}\quad\sum_{i=2}^{n}\zeta_{i}\stackrel{{\scriptstyle d}}{{\to}}N(0,1),$ (46)
-which establish
+which establish that
 $\xi_{n}=\sum_{i=2}^{n}\zeta_{i}-n^{-1/2}\frac{\widetilde{U}_{-1}^{\prime}P_{X_{2}(\rho_{n})}U}{\left(E(\widetilde{U}_{0}^{2}U_{1}^{2})\right)^{1/2}}\stackrel{{\scriptstyle d}}{{\to}}N(0,1).$ (47)
@@
 $=\frac{2h_{n}^{4}}{\frac{1}{6}h_{n}^{4}+O(h_{n}^{5})}+o(1)\to 12.$ (82)
-The proof is thus completed. ∎
+Thus, the proof is completed.. ∎
 Proof of Lemma 4.By setting $\phi_{n,i}=1$ in the proof of Lemma 5 in AG12, we have already established this lemma except for parts (e), (g), (i), (l), and certain combinations of part (n).
@@
 We provide additional simulation results in this section.
-Table SM.1 reports the CPs of [13, Mik07 hereafter] CIs.
+Table SM.1 reports the CPs of the [14, Mik07 hereafter] CI.
 [^4] The Mik07 CI differs from the [4] CI in that there is no initial-condition component in the asymptotic distribution of the test statistic, and a homoskedastic variance estimator is used.
-The CPs of Mik07 CIs are close to the nominal level when the $U_{i}$ sequences are i.i.d.
+The CPs of the Mik07 CI are close to the nominal level when the $U_{i}$ sequences are i.i.d.
 and $Y_{0}^{*}$ is either fixed or stationary.
-When $Y_{0}^{*}$ is drawn from scaled $n$ or explosive distributions, or when the $U_{i}$ sequences are non-i.i.d., the CPs of Mik07 CIs fall below 95%.
-Table SM.1: Coverage probabilities ($\times 100$) of nominal 95% Mik07 CIs
+When $Y_{0}^{*}$ is drawn from scaled $n$ or explosive distributions, or when the $U_{i}$ sequences are non-i.i.d., the CPs of the Mik07 CI fall below 95%.
+Table SM.1: Coverage probabilities ($\times 100$) of the nominal 95% Mik07 CI
 Initial Conditions: Fixed Stationary
@@
 and $Y_{0}^{*}$ is fixed or stationary.
-The ratios are between 1 and 1.12, suggesting that the ALs of ICR CIs are similar to those of Mik07 CIs when the true data generating process is such that the Mik07 inference method is valid.
-Tables SM.3 and SM.4 report the ALs of [4, AG14 hereafter] and Mik07 CIs, respectively.
+The ratios are between 1 and 1.12, suggesting that the ALs of the ICR CI are similar to those of the Mik07 CI when the true data generating process is such that the Mik07 inference method is valid.
+Tables SM.3 and SM.4 report the ALs of [4, AG14 hereafter] and the Mik07 CI, respectively.
 The results are consistent with the conclusions drawn from Tables 4 and SM.2.
-Table SM.2: Ratios of average lengths of nominal 95% ICR to Mik07 CIs
+Table SM.2: Ratios of the average lengths of the nominal 95% ICR to Mik07 CIs
 Initial Conditions: Fixed Stationary
@@
 i.i.d. 1.00 1.01 1.02 1.03 1.09 1.01 1.01 1.03 1.05 1.12
-Table SM.3: Average lengths of nominal 95% AG14 CIs
+Table SM.3: Average lengths of the nominal 95% AG14 CI
 Initial Conditions: Fixed Stationary
@@
 ARCH4 .35 .30 .24 .12 .03 .17 .15 .12 .05 .01
-Table SM.4: Average lengths of nominal 95% Mik07 CIs
+Table SM.4: Average lengths of the nominal 95% Mik07 CI
 Initial Conditions: Fixed Stationary